### 22 Jan 2024 v1

# A discrete model for Gell-Mann matrices

Abstract.
I propose a discrete model for the Gell-Mann matrices, which allows them to participate in discrete symmetries of three generations of four types of elementary fermions, in addition to their usual role in describing a continuous group $SU(3)$ of colour symmetries. This model sheds new light on the mathematical (rather than physical) necessity for ‘mixing’ between the various gauge groups $SU(3)$, $SU(2)$ and $U(1)$ of the Standard Model.
1. Introduction
1.1. Context, aims and objectives
The Gell-Mann matrices [1, 2] are a particular choice of orthonormal basis for the (complex) Lie algebra $\mathfrak{su}(3)$, analogous to the Pauli matrices which form a basis for $\mathfrak{su}(2)$. They are an essential part of the calculational tools of quantum chromodynamics (QCD) as a description of the strong nuclear force [3, 4]. The analogy with Pauli matrices is not complete, however, since the Pauli matrices are unitary, but the Gell-Mann matrices are not. In particular, the Pauli matrices are non-singular, so they can also be used to describe a finite group of discrete (unitary) symmetries. This is useful for describing the discrete symmetries of weak isospin, that distinguishes electrons from neutrinos.
Specifically, the Pauli matrices generate a commuting product of the cyclic group $Z_{4}$ of order $4$ with the quaternion group $Q_{8}$ of order $8$, giving a finite analogue of the gauge group $U(1)_{Y}\times SU(2)_{L}$ of weak hypercharge and weak isospin. But the ‘mixing’ of weak hypercharge and weak isospin to create electric charge implies that $U(1)_{em}$ does not commute with $SU(2)_{L}$. The only reasonable way to accommodate this property in the finite groups is to take a cyclic group $Z_{3}$ in place of $U(1)_{em}$, and let $Z_{3}$ act as automorphisms of $Q_{8}$. This leads to a group of order $24$, known as the binary tetrahedral group (among many other names), which is a semi-direct product $Q_{8}\rtimes Z_{3}$ of $Q_{8}$ and $Z_{3}$, rather than a direct product $Q_{8}\times Z_{3}$. Possible applications of this finite group to the modelling of elementary particles, and in particular of electro-weak mixing, are explored in [5, 6, 7, 8, 9, 10].
The Gell-Mann matrices, by contrast, are singular, so they do not generate a finite group, and they cannot be used for discrete symmetries such as the three-generation symmetry of electrons. This means that the three generations have to be added to the Standard Model ‘by hand’, rather than arising automatically from the formalism. The aim of this paper is to describe a method by which the Gell-Mann matrices can be given a discrete structure that might be useful for this purpose. I also explore options for mixing the strong force with the electro-weak forces by the process of using semi-direct products of the finite groups in place of the direct products of Lie groups. It turns out that there is a unique possibility, arising from a particular three-dimensional complex reflection group [11, 12] of order $648$.
The main aim of this paper is to explain the representation theory of this group of order $648$ in relation to the Standard Model. An important part of this aim is to understand how the action of $Z_{3}$ on $Q_{8}$ relates to the Standard Model implementation of electro-weak mixing. Similarly, the action of the Pauli matrices on the (modified) Gell-Mann matrices needs to be understood in the context of weak-strong mixing. In all cases, the finite group can give only the combinatorial structure of this mixing, without numerical values. The numerical values can only arise from choosing coordinates for the representations of the group. It is at this point that the finite groups are embedded in unitary gauge groups, $U(1)$, $SU(2)$ and $SU(3)$ respectively, and the structure of the Standard Model is recovered.
1.2. Pauli matrices
The standard textbook [2] Pauli matrices are
(7) $\sigma_{x}=\begin{pmatrix}0&1\cr 1&0\end{pmatrix},\quad\sigma_{y}=\begin{pmatrix}0&-i\cr i&0\end{pmatrix},\quad\sigma_{z}=\begin{pmatrix}1&0\cr 0&-1\end{pmatrix}$
that is, the Hermitian matrices in physicists’ convention. These give rise to elements of $SU(2)$ via complex exponentiation as follows:
(12) $\exp(i\sigma_{x}\theta)=\begin{pmatrix}\cos\theta&i\sin\theta\cr i\sin\theta&\cos\theta\end{pmatrix},\quad\exp(i\sigma_{y}\theta)=\begin{pmatrix}\cos\theta&\sin\theta\cr-\sin\theta&\cos\theta\end{pmatrix},$
(15) $\exp(i\sigma_{z}\theta)=\begin{pmatrix}e^{i\theta}&0\cr 0&e^{-i\theta}\end{pmatrix}.$
In addition there is a scalar matrix $i=\sigma_{x}\sigma_{y}\sigma_{z}$ which exponentiates to
(18) $\exp(i\theta)=\begin{pmatrix}e^{i\theta}&0\cr 0&e^{i\theta}\end{pmatrix}$
as elements of a scalar group $U(1)$. If this copy of $SU(2)$ is used for weak isospin, then the corresponding copy of $U(1)$ is used for weak hypercharge.
Mathematicians prefer to put the multiplication by $i$ into the definition of the matrices, instead of the exponentiation, so using the anti-Hermitian matrices
(25) $K=\begin{pmatrix}0&i\cr i&0\end{pmatrix},\quad J=\begin{pmatrix}0&1\cr-1&0\end{pmatrix},\quad I=\begin{pmatrix}i&0\cr 0&-i\end{pmatrix}$
in which the matrices $K,J,I$ behave like quaternions $IJ=-JI=K$ etc.
All the Pauli matrices, whether Hermitian or anti-Hermitian, are also unitary, as are the scalars. Therefore they generate finite subgroups of $U(2)$: in place of $U(1)_{Y}$ we have a cyclic group $Z_{4}$ of order $4$ generated by the scalar $i$, and in place of $SU(2)_{L}$ we have a quaternion group $Q_{8}$ of order $8$, generated by $i\sigma_{x}$, $i\sigma_{y}$ and $i\sigma_{z}$. In other words, there is a third interpretation of the Pauli matrices, as generators for unitary representations of a finite group, as well as the conventional interpretations as generators for unitary groups and Lie algebras. Such an interpretation is useful in cases (such as weak isospin) in which the underlying symmetry is discrete.
1.3. A $3$-dimensional analogue
The Gell-Mann matrices in physicists’ convention are $3\times 3$ analogues of the Hermitian traceless Pauli matrices:
(35) $\lambda^{1}=\begin{pmatrix}0&1&0\cr 1&0&0\cr 0&0&0\end{pmatrix},\quad\lambda^{2}=\begin{pmatrix}0&-i&0\cr i&0&0\cr 0&0&0\end{pmatrix},\quad\lambda^{3}=\begin{pmatrix}1&0&0\cr 0&-1&0\cr 0&0&0\end{pmatrix},$
(45) $\lambda^{4}=\begin{pmatrix}0&0&1\cr 0&0&0\cr 1&0&0\end{pmatrix},\quad\lambda^{5}=\begin{pmatrix}0&0&-i\cr 0&0&0\cr i&0&0\end{pmatrix},\quad\lambda^{6}=\begin{pmatrix}0&0&0\cr 0&0&1\cr 0&1&0\end{pmatrix},$
(52) $\lambda^{7}=\begin{pmatrix}0&0&0\cr 0&0&-i\cr 0&i&0\end{pmatrix},\quad\lambda^{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\cr 0&1&0\cr 0&0&-2\end{pmatrix}$
Again, the mathematicians’ convention is to multiply by $i$ to get anti-Hermitian matrices, which can then be exponentiated to get unitary matrices generating $SU(3)$. In both cases, the Gell-Mann matrices form a basis for the $8$-space of complex $3\times 3$ traceless matrices.
There are other possible bases, and it is suggested in [13] that a basis of real matrices might be more fundamental than Hermitian or anti-Hermitian matrices, for mathematical rather than physical reasons. But in none of these conventions are the matrices themselves unitary, so that the finite symmetries one might hope to see (such as the generation symmetry for fundamental fermions) are not directly available in the standard formalism. In fact it is easy to write down an orthonormal basis of unitary matrices, as I shall now demonstrate.
The natural analogue in three dimensions of the quaternion group (of order $82^{3}=$ 8) in two dimensions is the group $G_{27}$ of order $273^{3}=$ 27 generated by the matrices
(59) $\begin{pmatrix}1&0&0\cr 0&\exp(2\pi i/3)&0\cr 0&0&\exp(4\pi i/3)\end{pmatrix},\quad\begin{pmatrix}0&1&0\cr 0&0&1\cr 1&0&0\end{pmatrix}$
Writing $v=\exp(2\pi i/3)$ and $w=v^{2}=\exp(4\pi i/3)$ for simplicity, the trace zero matrices form a complex $8$-space with the following orthonormal basis:
(72) $\begin{pmatrix}1&0&0\cr 0&v&0\cr 0&0&w\end{pmatrix},\quad\begin{pmatrix}1&0&0\cr 0&w&0\cr 0&0&v\end{pmatrix},\quad\begin{pmatrix}0&1&0\cr 0&0&1\cr 1&0&0\end{pmatrix},\quad\begin{pmatrix}0&0&1\cr 1&0&0\cr 0&1&0\end{pmatrix},$
(85) $\begin{pmatrix}0&1&0\cr 0&0&v\cr w&0&0\end{pmatrix},\quad\begin{pmatrix}0&1&0\cr 0&0&w\cr v&0&0\end{pmatrix},\quad\begin{pmatrix}0&0&1\cr v&0&0\cr 0&w&0\end{pmatrix},\quad\begin{pmatrix}0&0&1\cr w&0&0\cr 0&v&0\end{pmatrix}.$
The first four are the generators (59) and their inverses, and the last four are the products of the first two with the next two. The generators (of order $3$) do not commute with each other, and the commutators $x^{-1}y^{-1}xy$ are scalar matrices of order $3$. In other words, there is a scalar group of order $3$ hidden in this group, consisting of scalars $1,v,w$, so that when we exponentiate these matrices there is a threefold ambiguity in the overall phase, corresponding to the fact that the Lie group $SU(3)$ has a centre consisting of scalars of order $3$, The above matrices are all unitary, and all have determinant $1$, so that the group of order $27$ that they generate is a finite subgroup of $SU(3)$.
Moreover, these $8$ matrices span the same complex $8$-space as the $8$ Gell-Mann matrices, namely the space of all complex matrices with trace zero. I therefore suggest that using these $8$ matrices in place of the $8$ Gell-Mann matrices makes essentially no difference to the Standard Model, as it is just a change of basis on the space of gluons. But it has the huge advantage over the Gell-Mann matrices, of incorporating finite (triplet) symmetries in a natural way. Indeed, it contains two independent triplet symmetries, and may therefore be able to accommodate generation symmetries as well as colour symmetries.
2. Extensions
2.1. The binary tetrahedral group
The quaternion group has an automorphism of order $3$, that cycles the elements $I$, $J$ and $K$, and that can be represented explicitly in the quaternion algebra by the matrix
(88) $W:=(-1+I+J+K)/2$ $=\frac{1}{2}\begin{pmatrix}-1+i&1+i\cr-1+i&-1-i\end{pmatrix}$
(91) $=\frac{1+i}{2}\begin{pmatrix}i&1\cr i&-1\end{pmatrix}$
This matrix extends the quaternion group $Q_{8}$ to a group of order $24$ that is known as the binary tetrahedral group. The above matrices give the ‘fundamental’ or ‘natural’ representation of this group inside $SU(2)$.
At the more abstract group theory level, we have constructed a semi-direct product of $Q_{8}$ by $Z_{3}$, which we denote $Q_{8}\rtimes Z_{3}$. The subgroup $Q_{8}$ generated by the matrices $K$, $J$, $I$ in (25) is normal, but the subgroup $Z_{3}$ generated by the above matrix $W$ is not. In place of the relations $IW=WI$, $JW=WJ$ that hold in the direct product, we have the relations $IW=WJ$ and $JW=WIJ$. The continuous analogue of $Q_{8}$ is $SU(2)$, and the continuous analogue of $Z_{3}$ is $U(1)$, so that the effect of this semi-direct product is to create a more complicated relationship between $SU(2)$ and $U(1)$ than the direct product $SU(2)\times U(1)$. In physics language this relationship is a ‘mixing’ of $SU(2)$ with $U(1)$ at the quantum level.
2.2. A ternary representation
In fact, the binary tetrahedral group can itself act as automorphisms of the group $G_{27}$ of order $27$ described in Section 1.3 above. We may first take generators for the quaternion subgroup $Q_{8}$ as
(98) $\frac{t}{3}\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix},\quad\frac{t}{3}\begin{pmatrix}1&v&v\cr w&v&w\cr w&w&v\end{pmatrix},$
where $t=v-w=\sqrt{-3}$, and check all the required relations explicitly. The calculations are all minor variants of the equation
(108) $\begin{pmatrix}0&1&0\cr 0&0&1\cr 1&0&0\end{pmatrix}\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix}$ $=\begin{pmatrix}1&v&w\cr 1&w&v\cr 1&1&1\end{pmatrix}$
(115) $=\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix}\begin{pmatrix}1&0&0\cr 0&v&0\cr 0&0&w\end{pmatrix}$
which shows how the first of the given automorphisms converts permutation matrices into diagonal matrices.
At this stage we have a group of order $216$ that combines finite analogues of $SU(3)$ and $SU(2)$ into a single group. This is a semi-direct product $G_{27}\rtimes Q_{8}$, in which the subgroup $G_{27}$ is normal, but the subgroup $Q_{8}$ is not. Since it is not a direct product, it does not give rise to a direct product $SU(3)\times SU(2)$, but instead ‘mixes’ the two gauge groups.
Finally we can add in a matrix with determinant $v$, such as
(119) $\begin{pmatrix}v&0&0\cr 0&1&0\cr 0&0&1\end{pmatrix}$
in order to incorporate $U(1)$ as well, extending the quaternion group to the binary tetrahedral group. This gives us altogether a group of order $6482^{3}.3^{4}=$ 648 (the triple cover of the Hessian group) that can be interpreted as a finite analogue of the complete gauge group $U(1)\times SU(2)\times SU(3)$ of the Standard Model.
But the finite group is not a direct product of three factors in the way that the Lie group must be. It is an iterated semi-direct product $Q_{27}\rtimes Q_{8}\rtimes Z_{3}$. Thus the finite analogues of the three factors are mixed together in quite a complicated way, which may have important consequences for the ‘mixing’ of the different forces in the Standard Model. We have already discussed electroweak mixing in the context of $Q_{8}\rtimes Z_{3}$, but here we have a much more complicated weak-strong mixing in $G_{27}\rtimes Q_{8}$, as well as electro-strong mixing in $G_{27}\rtimes Z_{3}$. Further details will be discussed below.
2.3. A complex reflection group
The group of order 648 is generated by the conjugates of (119), which has one eigenvalue $v$ and two eigenvalues $1$. An invertible matrix with a single non-trivial eigenvalue is known as a reflection. Real reflections have real eigenvalues, so the non-trivial eigenvalue is $-1$, but complex reflections can have any root of unity as an eigenvalue. Our group is therefore a $3$-dimensional complex reflection group. There are $24$ reflections altogether, in $12$ ‘mirrors’:
(123) $\begin{array}[]{cccc}(t,0,0)&(1,1,1)&(v,1,1)&(w,1,1)\cr(0,t,0)&(1,v,w)&(1,1,v)&(1,w,1)\cr(0,0,t)&(1,v,w)&(1,v,1)&1,1,w)\end{array}$
The two reflections in the mirror $r$ are obtained by putting $\lambda=v$ or $\lambda=w$ in:
(124) $x\mapsto x-\frac{x.r}{r.r}(1-\lambda)r$
The four mirrors given in the top row generate one of the nine copies of the binary tetrahedral group, acting on the $2$-dimensional subspace perpendicular to $(0,1,-1)$. In particular, this group is a $2$-dimensional complex reflection group. The images of the (anti-Hermitian) Pauli matrices can be written as products of two reflections of order $3$ (with opposite determinants). This representation is equivalent to the sum of the trivial representation and the usual representation on a complex $2$-space.
If we expand the complex $1$-spaces to real $2$-spaces, then the six complex scalar multiples of $12$ complex $3$-vectors become two real scalar multiples of $36$ real $6$-vectors, and the corresponding $36$ real reflections generate the Weyl group of $E_{6}$. In other words, the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ is a subgroup of the Weyl group of $E_{6}$, and is in fact a maximal subgroup of the rotation part of the Weyl group. It is possible that this fact may have led to the appearance of $E_{6}$ symmetry in certain Grand Unified Theories (GUTs) such as [14, 15].
3. Representations and characters
3.1. Continuous gauge groups
So far we have only looked at the combinatorial structure of the group, which gives a qualitative but not quantitative picture of the underlying physics. In order to introduce measurable physical quantities such as mass, momentum and energy, and allocate them to elementary (or composite) particles, we need to introduce real (or complex or quaternionic) representations, and use real (or complex or quaternionic) gauge groups to choose coordinates for the representations. There are three fundamental representations, described above, given by the embeddings of finite groups in Lie groups as follows:
(125) $Z_{3}$ $\subset U(1)$
(126) $Q_{8}\rtimes Z_{3}$ $\subset Sp(1)\cong SU(2)$
(127) $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ $\subset U(3)$
These embeddings show how the gauge groups $U(1)$, $SU(2)$ and $SU(3)$ of the Standard Model arise from the finite group. In each case, the real scalar factor has been taken out of the gauge group, since it only defines a unit of measurement, rather than a genuine physical property.
In the case of $U(1)$ acting on a complex $1$-dimensional representation, there is one physical parameter to measure, represented as an angle, or phase, in the complex plane. In the case of $SU(2)$ acting on a $1$-dimensional quaternionic representation, there are three such angles. Therefore by studying the representations of the binary tetrahedral group $Q_{8}\rtimes Z_{3}$, we might expect to find four of the fundamental mixing angles of the Standard Model, including the electro-weak mixing angle (Weinberg angle). A tentative identification of these four angles is obtained in [9].
In the case of $U(3)$, there are five independent angles in the $3$-dimensional complex representation, bringing the total number of mixing angles up to $9$, enough for the Weinberg angle plus four each in the CKM (Cabibbo–Kobayashi–Maskawa [16, 17]) and PMNS (Pontecorvo–Maki–Nakagawa–Sakata [18, 19]) matrices.
3.2. Clifford theory
The standard method of constructing representations and characters of semi-direct products is Clifford theory [20], which is also used more generally whenever a group has a normal subgroup. Since the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ has $6$ normal subgroups, forming a single chain, we can work our way down the chain, from the largest normal subgroup to the smallest, adding new representations at each stage. The method is standard, and the calculations straightforward, so I do not give many details.
At the first stage, we have only the trivial quotient group, and the trivial representation. At the second stage, we have a quotient $Z_{3}$, which adds a $2$-dimensional real representation, or equivalently a $1$-dimensional complex representation and its complex conjugate. At the third stage, the quotient group is the alternating group $A_{4}$ on four letters, and the new representation is the monomial representation of the group acting by conjugation on the three Pauli matrices, with signs attached.
At the fourth stage, we have the fundamental pseudoreal representation as quaternions, with $Z_{3}$ as the corresponding ‘inertial quotient’, so that the other representations at this stage are obtained by tensoring with the representations of $Z_{3}$. This gives us a pair of complex conjugate two-dimensional representations, or equivalently a real four-dimensional representation. Putting all these stages together gives us the representations of the binary tetrahedral group, which are very well known and have been widely studied. The character table can be found in [21, p.404] and is reproduced in Table 1 for convenience, in a shorthand form in which conjugacy classes that are scalar multiples of each other are grouped together:
Table 1. Character table of the binary tetrahedral group
±$\begin{array}[]{l|cccc}&\pm 1&K&\pm W&\pm W^{2}\cr\hline\cr\mathbf{1}a&1&1&1&1\cr\mathbf{1}b&1&1&v&w\cr\mathbf{1}c&1&1&w&v\cr\mathbf{3}&3&-1&0&0\cr\mathbf{2}a&\pm 2&0&\mp 1&\mp 1\cr\mathbf{2}b&\pm 2&0&\mp v&\mp w\cr\mathbf{2}c&\pm 2&0&\mp w&\mp v\cr\hline\cr\end{array}$
Of particular significance here is that the three faithful representations are obtained by taking the fundamental representation $\mathbf{2}a$ and tensoring with the representations $\mathbf{1}a$, $\mathbf{1}b$ and $\mathbf{1}c$ of $Z_{3}$. This type of pattern is a general feature of Clifford theory, and is a generalisation of the decomposition of representations of direct products of groups as tensor products of representations of the factors. In particular, we have a discrete version of the decomposition of representations of $U(1)\times SU(2)$, but with the ‘mixing’ of $U(1)$ with $SU(2)$ already incorporated into the mathematical structure. A combinatorial version of the mixing occurs already at the group-theoretical level, but a quantitative estimate of the mixing angle can only be obtained by looking in detail at the coordinates of the representations.
3.3. The Hessian group
Coming now to $G_{27}$, whose character table can be found in [21, p.400] we first take a $1$-dimensional complex representation, with inertial quotient $Z_{3}$, and induce up to the whole group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$, to get an $8$-dimensional real representation, together with its tensor products with the representations of $Z_{3}$. The scalars in $G_{27}$ act trivially on these representations, so they are representations of a group $(Z_{3}\times Z_{3})\rtimes Q_{8}\rtimes Z_{3}$ of order $216$, which has been known as the Hessian group since the 1870s.
The table of characters of the complex irreducible representations of the Hessian group is given in Table 2. The vertical lines delineate normal subgroups, so that the first two columns comprise the normal subgroup $G_{27}/Z_{3}\cong Z_{3}\times Z_{3}$, generated by the modified Gell-Mann matrices modulo scalars, and the next two columns comprise the quaternion group $Q_{8}$ generated by the Pauli matrices. The horizontal lines separate cohorts of characters for the various quotient groups. The top two rows contain the centralizer orders and the sizes of the conjugacy classes, respectively.
Table 2. The character table of the Hessian group
$\begin{array}[]{r|rr|rr|rrrrrr|}&216&27&24&4&18&9&6&18&9&6\cr&1&8&9&54&12&24&36&12&24&36\cr\hline\cr\mathbf{1}a&1&1&1&1&1&1&1&1&1&1\cr\mathbf{1}b&1&1&1&1&v&v&v&w&w&w\cr\mathbf{1}c&1&1&1&1&w&w&w&v&v&v\cr\hline\cr\mathbf{3}a&3&3&3&-1&0&0&0&0&0&0\cr\hline\cr\mathbf{2}a&2&2&-2&0&-1&-1&1&-1&-1&1\cr\mathbf{2}b&2&2&-2&0&-v&-v&v&-w&-w&w\cr\mathbf{2}c&2&2&-2&0&-w&-w&w&-v&-v&v\cr\hline\cr\mathbf{8}a&8&-1&0&0&2&-1&0&2&-1&0\cr\mathbf{8}b&8&-1&0&0&2v&-v&0&2w&-w&0\cr\mathbf{8}c&8&-1&0&0&2w&-w&0&2v&-v&0\cr\hline\cr\end{array}$
For some purposes it is useful to replace a pair of complex conjugate characters by their real sum, and to fuse a conjugacy class with its inverse class. This give us a simplified table (see Table 3).
Table 3. Real characters of the Hessian group
$\begin{array}[]{c|rr|rr|rrr|}&1&1&1&1&1&1&1\cr U(1)=SO(2)&2&2&2&2&-1&-1&-1\cr\hline\cr SO(3)&3&3&3&-1&0&0&0\cr\hline\cr SU(2)&2&2&-2&0&-1&-1&1\cr U(2)&4&4&-4&0&1&1&-1\cr\hline\cr SU(3)&8&-1&0&0&2&-1&0\cr U(3)&16&-2&0&0&-2&1&0\end{array}$
The first column gives the name of the compact Lie group generated by the finite group, from which we can see a close relationship to the gauge group of the Standard Model. The even-numbered rows contain the basic building blocks $U(1)$, $SU(2)$ and $SU(3)$ respectively, and the other rows contribute to the ‘mixing’ between these three groups.
3.4. The faithful characters
Finally we take a $3$-dimensional complex faithful representation of $G_{27}$, with inertial quotient $Q_{8}\rtimes Z_{3}$, so that Clifford theory implies that we must now tensor with all the representations of the latter group. Let us arrange the conjugacy classes grouped together into scalar multiples. Then we only need to specify the character value on one of these classes, and multiply by the appropriate scalars to get the others. In three cases an element is conjugate to its scalar multiples, so that the character values are all 00, while in the other seven they are not. The $14$ faithful complex characters are those given in Table 4 together with their complex conjugates.
Table 4. Abbreviated table of characters for quarks
$\begin{array}[]{r|rr|rr|rrrrrr|}\mathbf{3}b&3&0&-1&-1&t&0&t&-t&0&-t\cr\mathbf{3}c&3&0&-1&-1&vt&0&vt&-wt&0&-wt\cr\mathbf{3}d&3&0&-1&-1&wt&0&wt&-vt&0&-vt\cr\hline\cr\mathbf{9}&9&0&-3&1&0&0&0&0&0&0\cr\hline\cr\mathbf{6}a&6&0&2&0&-t&0&t&t&0&-t\cr\mathbf{6}b&6&0&2&0&-vt&0&vt&wt&0&-wt\cr\mathbf{6}c&6&0&2&0&-wt&0&wt&vt&0&-vt\end{array}$
It is not really necessary to write out this table in full, since, as noted above, the seven rows are obtained by multiplying the first row with the seven irreducible characters of the binary tetrahedral group, that is the first seven characters of Table 2. If we take the latter to describe properties of leptons, then Table 4 describes the corresponding properties of quarks. Indeed, if we have the representations $\mathbf{1}b$, $\mathbf{2}a$ and $\mathbf{3}b$ then we can calculate all the other representations easily as tensor products, subject only to the relations
(128) $\mathbf{2}a\otimes\mathbf{2}a$ $=\mathbf{1}a+\mathbf{3}a$
(129) $\mathbf{3}b\otimes\overline{\mathbf{3}b}$ $=\mathbf{1}a+\mathbf{8}a.$
Hence it is only necessary to gauge the representations $\mathbf{1}a$, $\mathbf{2}a$ and $\mathbf{3}b$, using the gauge groups $U(1)$, $SU(2)$ and $SU(3)$ respectively.
4. Computing gauges
4.1. The electroweak gauge
The $U(1)$ gauge group contains only one parameter, or ‘mixing angle’, that identifies the directions of the real and imaginary parts of the representation. This angle has been computed in [9] as $33.024^{\circ}$. The method of calculation was to superimpose the mass axis on an equilateral triangle marked with the three generations of electrons on the vertices. Hence it is expected that this mixing angle is one of the lepton mixing angles in the Standard Model.
The $SU(2)$ gauge group similarly contains exactly three independent parameters. The angles calculated in [9] were $28.165516^{\circ}$, $66.7276^{\circ}$ and $2.337325^{\circ}$. The method here was to identify discrete quaternions appropriate for four particles (the three generations of electron, plus a proton) that are ‘fundamental’ as far as electroweak interactions are concerned, and again superimpose a mass axis. We should certainly expect one of these angles to be the electroweak mixing angle, but it is not clear a priori whether the other two are lepton-mixing (PMNS) or quark-mixing (CKM) angles. However, if we take into account that the embedding of $Q_{8}\rtimes Z_{3}$ into $SU(3)$ is described by the representation $\mathbf{1}b+\mathbf{2}c$, we should expect one ‘scalar’ (i.e. CP-violating phase) and one generation-mixing angle.
It is therefore proposed to identify these angles, calculated from the directions of the mass axis in the two cases, with four of the Standard Model mixing angles, as follows:
(130) $\theta_{12}^{PMNS}$ $={33.41^{\circ}}^{+.75^{\circ}}_{-.72^{\circ}}$
(131) $\theta_{W}$ $=28.172^{\circ}\pm.021^{\circ}$
(132) $\delta_{CP}^{CKM}$ $=68.8^{\circ}\pm 4.5^{\circ}$
(133) $\theta_{23}^{CKM}$ $=2.38^{\circ}\pm.06^{\circ}.$
4.2. The strong gauge
The $SU(3)$ gauge group acts on a unitary $3$-space, which in principle requires $8$ ‘mixing angles’ to define a unitary basis. But we have already identified $3$ of them from an embedding of $SU(2)$ in $SU(3)$ (the fourth mixing angle above requires an extension to $U(3)$). Hence there are just five more mixing angles to identify.
In the Standard Model these are
(134) $\theta_{12}^{CKM}$ $=13.04^{\circ}\pm.05^{\circ},$
(135) $\theta_{13}^{CKM}$ =.$201^{\circ}\pm.011^{\circ},= .$
(136) $\theta_{13}^{PMNS}$ $={8.54^{\circ}}^{+.11^{\circ}}_{-.12^{\circ}},$
(137) $\theta_{23}^{PMNS}$ $={49.1^{\circ}}^{+1.0^{\circ}}_{-1.3^{\circ}},$
(138) $\delta_{CP}^{PMNS}$ $={197^{\circ}}^{+42^{\circ}}_{-25^{\circ}}.$
In principle it should be possible to calculate these from an appropriate set of particle masses, but it is far from obvious how many masses we need, or what the formulae are. Certainly we need at least six masses, which must presumably include all the quark masses, but it seems more likely that we need to include the three electron masses and the proton mass as well, making ten in all.
4.3. Identifying particles
However it is done, we must start by identifying the particles we use with certain vectors or subspaces in $\mathbb{C}^{3}$. Therefore we must begin by examining the geometry of this $3$-dimensional space, using the complex reflection group as a guide. The most important vectors are those that lie in the $12$ ‘mirrors’. The particular scalar multiple chosen is irrelevant for defining the reflection, so to begin with at least we can ignore the scalar multiplication. The Hessian group then acts on this set of $12$ one-spaces in four triplets, as shown in (123). The reflections in any mirror fix the mirrors in its column, and cycle the other three columns. The normal subgroup $G_{27}/Z_{3}$ of order $9$ contains four subgroups of order $3$, each of which fixes the mirrors in one column, and cycles the mirrors in the other three columns (not always in the same direction!). It is therefore tempting to identify these $12$ mirrors with the $12$ types of elementary fermions, and use the scalars for unobservable properties such as colours of quarks and spin direction of electrons. However, the details of such an identification are tricky.
An alternative approach might be to fix a particular mirror, say $(t,0,0)$, so that the reflections rotate the mirror itself, to $(vt,0,0)$ and $(wt,0,0)$, fix the two perpendicular mirrors $(0,t,0)$ and $(0,0,t)$, and move the other $9$ mirrors. We can then perhaps identify the last $9$ mirrors as three generations of electrons and up and down quarks in some way, in order to get masses into the picture. Then we might use the mirror itself for a tenth mass such as a proton, or alternatively treat one or all of the three mirrors in the first column as neutrinos. The latter suggestion appears to mix the electron generation with the spin direction for a neutrino, which is exactly the kind of mathematical property that is needed both for the chirality of the weak interaction detected by the Wu experiment [22] and for explaining neutrino oscillations [23].
There is another set of $9$ one-spaces defined as the fixed spaces of the $9$ copies of $Q_{8}\rtimes Z_{3}$. These are the images of $(0,1,-1)$ under coordinate permutations and scalar multiplications on the coordinates. If we fix a particular copy of the group, then the $9$ one-spaces split as $1+8$, which suggests a splitting of the massless bosons in the Standard Model into one photon plus eight gluons. If we have identified the individual coordinates with neutrinos as suggested above, then there is a possible interpretation of gluons (and the photon) as bound states of a neutrino and an anti-neutrino. Indeed, if we take the neutrinos as a basis for $\mathbf{3}b$ and the anti-neutrinos as a basis for $\overline{\mathbf{3}b}$, as above, then the bound states should lie in $\mathbf{3}b\otimes\overline{\mathbf{3}b}=\mathbf{1}a+\mathbf{8}a$.
5. Geometry
5.1. Finite fields
The mathematical structure is best elucidated by regarding $1,v,w$ not as complex numbers, but as abstract elements of the field of order $4$. This is a field of characteristc $2$, so that $2=0$, $-1=1$ and $3=1$. The equation $01+v+w=$ 0 still holds, as in the complex numbers, so that $1+v=w$ and $1+w=v$, as well as $v+w=1$. The multiplicative rules $v^{2}=w$, $w^{2}=v$ and $vw=1$ also hold just as before, and the involution that swaps $v$ with $w$ is the appropriate analogue of complex conjugation. Denote this field by $\mathbb{F}_{4}$, and denote the $n$-dimensional unitary group by $U(n,\mathbb{F}_{4})$.
Then it is easy to see that $U(1,\mathbb{F}_{4})$ is the cyclic group of order $3$, consisting of all non-zero $1\times 1$ matrices. The whole group of order $648$ constructed above is then $U(3,\mathbb{F}_{4})$. But the quaternion group and the binary tetrahedral group do not have obvious names in terms of $\mathbb{F}_{4}$.
The reason for this is that the Pauli matrices are more naturally written over the field $\mathbb{F}_{9}$ of order $9$, whose non-zero elements are ±$\pm 1$, ±$\pm i$ and ±$\pm 1\pm i$. We can then see that the binary tetrahedral group is isomorphic to $SU(2,\mathbb{F}_{9})$. From this point of view, therefore, the reason for the complicated mixing between $SU(2)$ and $SU(3)$ in the Standard Model is ultimately a complicated relationship between the finite fields of orders $4$ and $9$. The particular feature that makes the finite groups behave completely differently from Lie groups is the chain of quotient maps
(139) $U(3,\mathbb{F}_{4})\rightarrow SU(2,\mathbb{F}_{9})\rightarrow U(1,\mathbb{F}_{4})$
This provides a unification of $U(3)$ and $SU(2)$ that is specific to these finite fields, and occurs nowhere else in mathematics. In particular, such unification cannot be achieved with Lie groups over the complex numbers.
The $3$-dimensional space over $\mathbb{F}_{4}$ on which $U(3,\mathbb{F}_{4})$ acts contains $634^{3}-1=$ 63 vectors, each in a set of $3$ scalar multiples, and therefore forming $63/3=21$ one-dimensional subspaces. These $21$ ‘points’ of the geometry divide into $12$ ’odd’ (non-singular) and $9$ ’even’ (singular) points. Each point is perpendicular to a ‘line’. The $12$ non-singular lines each contain $2$ non-singular and $3$ singular points, while the $9$ singular lines each contain $1$ singular point and $4$ non-singular points.
The points can be arranged as
(146) $\begin{array}[]{|cccc|}100&111&v11&w11\cr 010&1vw&11v&1w1\cr 001&1wv&1v1&11w\end{array}\qquad\begin{array}[]{|ccc|}011&101&110\cr 0vw&w0v&vw0\cr 0wv&v0w&wv0\end{array}$
to exhibit the symmetry. These are just the reductions to $\mathbb{F}_{4}$ of the $12$ ‘mirrors’ (corresponding to $Z_{3}$ subgroups) and $9$ copies of the Pauli matrices (corresponding to $Q_{8}$ subgroups), so that all of the above discussion of particles applies equally to the finite geometry.
5.2. Interpretations and speculations
In particular, we can interpret the $12$ non-singular points as the fundamental fermions, arranged into four columns of three generations (flavours) each, and it would seem to make sense to interpret the first two columns as leptons, and the last two as quarks, although the mathematics allows us an arbitrary choice here. It is noticeable, however, that in this interpretation there is no single ‘generation symmetry’ that acts on all four columns simultaneously. Instead, there is a group of order $9$ containing elements of order $3$ that act on any three of the four columns.
This is physically reasonable, since it is known that quark generations behave differently from lepton generations in the weak interaction, so that at least two different generation symmetries are required in order to describe what is experimentally observed. We can choose the two generation symmetries in a variety of ways, for example an electron generation symmetry that acts on electrons but not neutrinos, and a neutrino generation symmetry that acts on neutrinos but not electrons. Such a choice shows explicitly that the neutrinos can change flavour independently of the electrons, and therefore models the observed behaviour of neutrino oscillations between the three flavours.
In other words, the proposed splitting of the ‘generation’ symmetry into two separate triplet symmetries is strongly supported by experiment. But there is a crucial difference between the electron generation symmetry, in which the mass changes very obviously by several orders of magnitude between generations, and the neutrino generation symmetry, in which no mass can be detected directly at all, but is instead inferred from the details of the theoretical model.
It is reasonable to suppose, therefore, that the electron generation symmetry acts to change the mass of the quarks, while the neutrino generation symmetry has no detectable effect on quark mass. In other words, the neutrino generation symmetry effectively acts on the colours of the quarks rather than their generation. But this implies that the neutrino generation should rather be interpreted as a colour, and that neutrino oscillation should be interpreted as a change in colour, and must therefore be mediated by gluons.
This radical proposal has a number of interesting consequences. First, it allows neutrinos to change generation/colour by an interaction (with a massless gluon), rather than spontaneously, and therefore it allows this to happen even if the neutrinos have exactly zero mass. It therefore does not require a hypothesis of non-zero neutrino mass. Second, it allows colours to escape from hadrons, attached to neutrinos, and therefore removes the need to explain ‘colour confinement’. Third, it allows free massless gluons to exist, which are then available as mediators for a quantum theory of gravity. Indeed, it is possible that it even allows bound states of two gluons (with different colour charges) to form (massless) spin $2$ gravitons.
6. Conclusion
6.1. Summary
The suggestion that physics might be fundamentally discrete, and that the apparently continuous behaviour we often observe is an illusion, goes back a long way. In the context of quantum mechanics, this suggestion was made by Einstein [24] as early as 1935, but no discrete theory was found to rival the continuous models based on the Schrödinger equation and then the Dirac equation. The Standard Model of Particle Physics, largely completed by the mid 1970s, is entirely based on continuous groups, so much so that the emergence of discrete experimental observations is still not explained in a totally convincing way [25, 26, 27].
Theories beyond the Standard Model are almost without exception continuous theories. A few lone voices, such as ’t Hooft [28], still speak out in favour of Einstein’s dream of a discrete foundation, even if this has to be at the Planck scale. But a convincing discrete model is still lacking, and few people are seriously looking for one. The ones I have put forward in [9, 29, 30, 31, 10, 32] all suffer from drawbacks of various kinds, most notably the lack of any convincing analogue of the colour $SU(3)$, which is an essential part of the Standard Model.
Nevertheless, it is a mathematical fact that finite groups are much more complicated than Lie groups, and quite small finite groups can model properties that cannot be modelled with Lie groups at all. In this paper, therefore, I have tried to remedy this defect by explicitly proposing a finite analogue or precursor of $SU(3)$. At a combinatorial (qualitative) level it has many desirable properties that mirror known experimental properties of elementary particles, in particular the chirality of the weak interaction, and the mixing of electromagnetism and the weak and strong nuclear forces. Quantitative analysis of masses, mixing angles and probability amplitudes is beyond the scope of this paper.
6.2. Further work
The proposed unification of Gell-Mann and Pauli matrices using a 3-dimensional complex reflection group really does nothing more than provide a discrete foundation for the Standard Model of Particle Physics as it currently stands. Nevertheless, it permits new methods of calculation, using finite geometries, which have been shown to be effective in calculating four of the mixing angles from fundamental mass parameters. A priority for further work is therefore to try to find similar calculations for the remaining five mixing angles.
Further unification with the Dirac matrices cannot be done in the Standard Model, since it is impossible to arrange for the Lorentz group to commute with the whole of the gauge group in this case. What this means is that a putative unification with Dirac matrices provides a possible mechanism for particle physics to influence the shape of spacetime, and hence couple to gravity. Investigating this possibility further is a priority for future work.
If a quantum theory of gravity along these lines turns out to be viable, then it might lead to methods for calculating the $15$ mass parameters of the Standard Model from something more fundamental. It is not at all clear what could be ‘more fundamental’, or how it could be used, so that this is at present a highly speculative idea. Nevertheless, it is clear that it will be necessary to examine closely the theoretical and practical definitions of inertial frames of reference in different contexts, and to calculate the effects of transformations between them.
In order to gauge $15$ masses we need an $8$-dimensional complex representation, or a $4$-dimensional quaternionic representation. Since $SL(4,\mathbb{H})$ is a real form of $SU(8)$, of Lie type $A_{7}$, and since $A_{7}$ is contained in both $E_{7}$ and $D_{8}$, there are a number of ways to embed such a gauge group in (complex) $E_{8}$. It may even be possible to embed $SL(4,\mathbb{H})$ in a real form of $E_{8}$. In the $E_{7}$ case, however, there are no spinor representations of $SL(4,\mathbb{H})$ in $E_{8}$, which seems to rule this one out. In the $D_{8}$ case, the spinors split into $4+28$ quaternionic dimensions, which may be viable as a splitting into neutrinos versus the charged fundamental fermions. Even so, the model would still be quite unlike the various $E_{8}$ models [13, 33, 34, 35] that have been proposed over a number of years.