### 25 Jan 2024 v2

# A discrete model for Gell-Mann matrices

Abstract.
I propose a discrete model for the Gell-Mann matrices, which allows them to participate in discrete symmetries of three generations of four types of elementary fermions, in addition to their usual role in describing a continuous group $SU(3)$ of colour symmetries. This model sheds new light on the mathematical (rather than physical) necessity for ‘mixing’ between the various gauge groups $SU(3)$, $SU(2)$ and $U(1)$ of the Standard Model.
1. Introduction
1.1. Context, aims and objectives
The Gell-Mann matrices [1, 2] are a particular choice of orthonormal basis for the (complex) Lie algebra $\mathfrak{su}(3)$, analogous to the Pauli matrices which form a basis for $\mathfrak{su}(2)$. They are an essential part of the calculational tools of quantum chromodynamics (QCD) as a description of the strong nuclear force [3, 4]. The analogy with Pauli matrices is not complete, however, since the Pauli matrices are unitary, but the Gell-Mann matrices are not. In particular, the Pauli matrices are non-singular, so they can also be used to describe a finite group of discrete (unitary) symmetries. This is useful for describing the discrete symmetries of weak isospin, that distinguishes electrons from neutrinos.
Specifically, the Pauli matrices generate a commuting product of the cyclic group $Z_{4}$ of order $4$ with the quaternion group $Q_{8}$ of order $8$, giving a finite analogue of the gauge group $U(1)_{Y}\times SU(2)_{L}$ of weak hypercharge and weak isospin. But the ‘mixing’ of weak hypercharge and weak isospin to create electric charge implies that $U(1)_{em}$ does not commute with $SU(2)_{L}$. The only reasonable way to accommodate this property in the finite groups is to take a cyclic group $Z_{3}$ in place of $U(1)_{em}$, and let $Z_{3}$ act as automorphisms of $Q_{8}$. This leads to a group of order $24$, known as the binary tetrahedral group (among many other names), which is a semi-direct product $Q_{8}\rtimes Z_{3}$ of $Q_{8}$ and $Z_{3}$, rather than a direct product $Q_{8}\times Z_{3}$. Possible applications of this finite group to the modelling of elementary particles, and in particular of electro-weak mixing, are explored in [5, 6, 7, 8, 9, 10].
The Gell-Mann matrices, by contrast, are singular, so they do not generate a finite group, and they cannot be used for discrete symmetries such as the three-generation symmetry of electrons. This means that the three generations have to be added to the Standard Model ‘by hand’, rather than arising automatically from the formalism. The aim of this paper is to describe a method by which the Gell-Mann matrices can be given a discrete structure that might be useful for this purpose. I also explore options for mixing the strong force with the electro-weak forces by the process of using semi-direct products of the finite groups in place of the direct products of Lie groups. It turns out that there is a unique possibility, arising from a particular three-dimensional complex reflection group [11, 12] of order $648$.
The main aim of this paper is to explain the representation theory of this group of order $648$ in relation to the Standard Model. An important part of this aim is to understand how the action of $Z_{3}$ on $Q_{8}$ relates to the Standard Model implementation of electro-weak mixing. Similarly, the action of the Pauli matrices on the (modified) Gell-Mann matrices needs to be understood in the context of weak-strong mixing. In all cases, the finite group can give only the combinatorial structure of this mixing, without numerical values. The numerical values can only arise from choosing coordinates for the representations of the group. It is at this point that the finite groups are embedded in unitary gauge groups, $U(1)$, $SU(2)$ and $SU(3)$ respectively, and the structure of the Standard Model is recovered.
A secondary aim is to show how the (modified) Gell-Mann matrices can act on an extended (quaternionic) version of the Dirac matrices, and thereby change the local shape of spacetime. We then investigate (at a qualitative level) the possibility of using this action to develop a generally covariant model of quantum gravity, that appears to reduce to General Relativity [13, 14, 15] if the three generations of electrons are ignored, and all matter is assumed to consist of neutrons.
1.2. Pauli matrices
The standard textbook [2] Pauli matrices are
(7) $\sigma_{x}=\begin{pmatrix}0&1\cr 1&0\end{pmatrix},\quad\sigma_{y}=\begin{pmatrix}0&-i\cr i&0\end{pmatrix},\quad\sigma_{z}=\begin{pmatrix}1&0\cr 0&-1\end{pmatrix}$
that is, the Hermitian matrices in physicists’ convention. These give rise to elements of $SU(2)$ via complex exponentiation as follows:
(12) $\exp(i\sigma_{x}\theta)=\begin{pmatrix}\cos\theta&i\sin\theta\cr i\sin\theta&\cos\theta\end{pmatrix},\quad\exp(i\sigma_{y}\theta)=\begin{pmatrix}\cos\theta&\sin\theta\cr-\sin\theta&\cos\theta\end{pmatrix},$
(15) $\exp(i\sigma_{z}\theta)=\begin{pmatrix}e^{i\theta}&0\cr 0&e^{-i\theta}\end{pmatrix}.$
In addition there is a scalar matrix $i=\sigma_{x}\sigma_{y}\sigma_{z}$ which exponentiates to
(18) $\exp(i\theta)=\begin{pmatrix}e^{i\theta}&0\cr 0&e^{i\theta}\end{pmatrix}$
as elements of a scalar group $U(1)$. If this copy of $SU(2)$ is used for weak isospin, then the corresponding copy of $U(1)$ is used for weak hypercharge.
Mathematicians prefer to put the multiplication by $i$ into the definition of the matrices, instead of the exponentiation, so using the anti-Hermitian matrices
(25) $K=\begin{pmatrix}0&i\cr i&0\end{pmatrix},\quad J=\begin{pmatrix}0&1\cr-1&0\end{pmatrix},\quad I=\begin{pmatrix}i&0\cr 0&-i\end{pmatrix}$
in which the matrices $K,J,I$ behave like quaternions $IJ=-JI=K$ etc.
All the Pauli matrices, whether Hermitian or anti-Hermitian, are also unitary, as are the scalars. Therefore they generate finite subgroups of $U(2)$: in place of $U(1)_{Y}$ we have a cyclic group $Z_{4}$ of order $4$ generated by the scalar $i$, and in place of $SU(2)_{L}$ we have a quaternion group $Q_{8}$ of order $8$, generated by $i\sigma_{x}$, $i\sigma_{y}$ and $i\sigma_{z}$. In other words, there is a third interpretation of the Pauli matrices, as generators for unitary representations of a finite group, as well as the conventional interpretations as generators for unitary groups and Lie algebras. Such an interpretation is useful in cases (such as weak isospin) in which the underlying symmetry is discrete.
1.3. A $3$-dimensional analogue
The Gell-Mann matrices in physicists’ convention are $3\times 3$ analogues of the Hermitian traceless Pauli matrices:
(35) $\lambda^{1}=\begin{pmatrix}0&1&0\cr 1&0&0\cr 0&0&0\end{pmatrix},\quad\lambda^{2}=\begin{pmatrix}0&-i&0\cr i&0&0\cr 0&0&0\end{pmatrix},\quad\lambda^{3}=\begin{pmatrix}1&0&0\cr 0&-1&0\cr 0&0&0\end{pmatrix},$
(45) $\lambda^{4}=\begin{pmatrix}0&0&1\cr 0&0&0\cr 1&0&0\end{pmatrix},\quad\lambda^{5}=\begin{pmatrix}0&0&-i\cr 0&0&0\cr i&0&0\end{pmatrix},\quad\lambda^{6}=\begin{pmatrix}0&0&0\cr 0&0&1\cr 0&1&0\end{pmatrix},$
(52) $\lambda^{7}=\begin{pmatrix}0&0&0\cr 0&0&-i\cr 0&i&0\end{pmatrix},\quad\lambda^{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\cr 0&1&0\cr 0&0&-2\end{pmatrix}.$
Again, the mathematicians’ convention is to multiply by $i$ to get anti-Hermitian matrices, which can then be exponentiated to get unitary matrices generating $SU(3)$. In both cases, the Gell-Mann matrices form a basis for the $8$-space of complex $3\times 3$ traceless matrices.
There are other possible bases, and it is suggested in [16] that a basis of real matrices might be more fundamental than Hermitian or anti-Hermitian matrices, for mathematical rather than physical reasons. But in none of these conventions are the matrices themselves unitary, so that the finite symmetries one might hope to see (such as the generation symmetry for fundamental fermions) are not directly available in the standard formalism. In fact it is easy to write down an orthonormal basis of unitary matrices, as I shall now demonstrate.
The natural analogue in three dimensions of the quaternion group (of order $82^{3}=$ 8) in two dimensions is the group $G_{27}$ of order $273^{3}=$ 27 generated by the matrices
(59) $\begin{pmatrix}1&0&0\cr 0&\exp(2\pi i/3)&0\cr 0&0&\exp(4\pi i/3)\end{pmatrix},\quad\begin{pmatrix}0&1&0\cr 0&0&1\cr 1&0&0\end{pmatrix}$
Writing $v=\exp(2\pi i/3)$ and $w=v^{2}=\exp(4\pi i/3)$ for simplicity, the trace zero matrices form a complex $8$-space with the following orthonormal basis:
(72) $\begin{pmatrix}1&0&0\cr 0&v&0\cr 0&0&w\end{pmatrix},\quad\begin{pmatrix}1&0&0\cr 0&w&0\cr 0&0&v\end{pmatrix},\quad\begin{pmatrix}0&1&0\cr 0&0&1\cr 1&0&0\end{pmatrix},\quad\begin{pmatrix}0&0&1\cr 1&0&0\cr 0&1&0\end{pmatrix},$
(85) $\begin{pmatrix}0&1&0\cr 0&0&v\cr w&0&0\end{pmatrix},\quad\begin{pmatrix}0&1&0\cr 0&0&w\cr v&0&0\end{pmatrix},\quad\begin{pmatrix}0&0&1\cr v&0&0\cr 0&w&0\end{pmatrix},\quad\begin{pmatrix}0&0&1\cr w&0&0\cr 0&v&0\end{pmatrix}.$
The first four are the generators (59) and their inverses, and the last four are the products of the first two with the next two. The generators (of order $3$) do not commute with each other, and the commutators $x^{-1}y^{-1}xy$ are scalar matrices of order $3$. In other words, there is a scalar group of order $3$ hidden in this group, consisting of scalars $1,v,w$, so that when we exponentiate these matrices there is a threefold ambiguity in the overall phase, corresponding to the fact that the Lie group $SU(3)$ has a centre consisting of scalars of order $3$, The above matrices are all unitary, and all have determinant $1$, so that the group of order $27$ that they generate is a finite subgroup of $SU(3)$.
Moreover, these $8$ matrices span the same complex $8$-space as the $8$ Gell-Mann matrices, namely the space of all complex matrices with trace zero. I therefore suggest that using these $8$ matrices in place of the $8$ Gell-Mann matrices makes essentially no difference to the Standard Model, as it is just a change of basis on the space of gluons. But it has the huge advantage over the Gell-Mann matrices, of incorporating finite (triplet) symmetries in a natural way. Indeed, it contains two independent triplet symmetries, and may therefore be able to accommodate generation symmetries as well as colour symmetries.
2. Extensions
2.1. The binary tetrahedral group
The quaternion group has an automorphism of order $3$, that cycles the elements $I$, $J$ and $K$, and that can be represented explicitly in the quaternion algebra by the matrix
(88) $W:=(-1+I+J+K)/2$ $=\frac{1}{2}\begin{pmatrix}-1+i&1+i\cr-1+i&-1-i\end{pmatrix}$
(91) $=\frac{1+i}{2}\begin{pmatrix}i&1\cr i&-1\end{pmatrix}$
This matrix extends the quaternion group $Q_{8}$ to a group of order $24$ that is known as the binary tetrahedral group. The above matrices give the ‘fundamental’ or ‘natural’ representation of this group inside $SU(2)$.
At the more abstract group theory level, we have constructed a semi-direct product of $Q_{8}$ by $Z_{3}$, which we denote $Q_{8}\rtimes Z_{3}$. The subgroup $Q_{8}$ generated by the matrices $K$, $J$, $I$ in (25) is normal, but the subgroup $Z_{3}$ generated by the above matrix $W$ is not. In place of the relations
(92) $IZ=ZI,\quad JZ=ZJ$
that define the direct product, $Q_{8}\times Z_{3}=\langle I,J\rangle\times\langle Z\rangle$, we have the relations
(93) $IW=WJ,\quad JW=WIJ$
in the semidirect product $Q_{8}\rtimes Z_{3}=\langle I,J\rangle\rtimes\langle W\rangle$.
The continuous analogue of $Q_{8}$ is $SU(2)$, and the continuous analogue of $Z_{3}$ is $U(1)$, so that the effect of this semi-direct product is to create a more complicated relationship between $SU(2)$ and $U(1)$ than the direct product $SU(2)\times U(1)$. In physics language this relationship is a ‘mixing’ of $SU(2)$ with $U(1)$ at the quantum level. In the case of electro-weak mixing, $Z$ represents a generator for the group $U(1)_{Y}$ of weak hypercharge, while $W$ represents a generator of the group $U(1)_{em}$ of (electromagnetic) charge, and $I,J,K$ represent the three components of weak isospin, as generators for $SU(2)_{L}$.
2.2. A ternary representation
In fact, the binary tetrahedral group can itself act as automorphisms of the group $G_{27}$ of order $27$ described in Section 1.3 above. We may first take generators for the quaternion subgroup $Q_{8}$ as
(100) $-\frac{t}{3}\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix},\quad-\frac{t}{3}\begin{pmatrix}1&v&v\cr w&v&w\cr w&w&v\end{pmatrix},$
where $t=v-w=\sqrt{-3}$, and check all the required relations explicitly. The calculations are all minor variants of the equation
(110) $\begin{pmatrix}0&1&0\cr 0&0&1\cr 1&0&0\end{pmatrix}\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix}$ $=\begin{pmatrix}1&v&w\cr 1&w&v\cr 1&1&1\end{pmatrix}$
(117) $=\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix}\begin{pmatrix}1&0&0\cr 0&v&0\cr 0&0&w\end{pmatrix}$
which shows how the first of the given automorphisms converts permutation matrices into diagonal matrices.
At this stage we have a group of order $216$ that combines finite analogues of $SU(3)$ and $SU(2)$ into a single group. This is a semi-direct product $G_{27}\rtimes Q_{8}$, in which the subgroup $G_{27}$ is normal, but the subgroup $Q_{8}$ is not. Since it is not a direct product, it does not give rise to a direct product $SU(3)\times SU(2)$, but instead ‘mixes’ the two gauge groups. If we compute the action of the matrices (100) on the 3-vectors, we find that $(0,1,-1)$ is an eigenvector, with eigenvalue $1$. Hence $Q_{8}$ embeds in a subgroup $SU(2)$ of $SU(3)$, which gives the required mixing.
Finally we can add in a matrix with determinant $v$, such as
(121) $\begin{pmatrix}v&0&0\cr 0&1&0\cr 0&0&1\end{pmatrix}$
in order to incorporate $U(1)$ as well, extending the quaternion group to the binary tetrahedral group. This gives us altogether a group of order $6482^{3}.3^{4}=$ 648 (the triple cover of the Hessian group) that can be interpreted as a finite analogue of the complete gauge group $U(1)\times SU(2)\times SU(3)$ of the Standard Model.
But the finite group is not a direct product of three factors in the way that the Lie group must be. It is an iterated semi-direct product $Q_{27}\rtimes Q_{8}\rtimes Z_{3}$. Thus the finite analogues of the three factors are mixed together in quite a complicated way, which may have important consequences for the ‘mixing’ of the different forces in the Standard Model. We have already discussed electroweak mixing in the context of $Q_{8}\rtimes Z_{3}$, but here we have a much more complicated weak-strong mixing in $G_{27}\rtimes Q_{8}$, as well as electro-strong mixing in $G_{27}\rtimes Z_{3}$. Further details will be discussed below.
2.3. A complex reflection group
The group of order 648 is generated by the conjugates of (121), which has one eigenvalue $v$ and two eigenvalues $1$. An invertible matrix with a single non-trivial eigenvalue is known as a reflection. Real reflections have real eigenvalues, so the non-trivial eigenvalue is $-1$, but complex reflections can have any root of unity as an eigenvalue. Our group is therefore a $3$-dimensional complex reflection group. There are $24$ reflections altogether, in $12$ ‘mirrors’:
(125) $\begin{array}[]{cccc}(t,0,0)&(1,1,1)&(v,1,1)&(w,1,1)\cr(0,t,0)&(1,v,w)&(1,1,v)&(1,w,1)\cr(0,0,t)&(1,v,w)&(1,v,1)&1,1,w)\end{array}$
The two reflections in the mirror $r$ are obtained by putting $\lambda=v$ or $\lambda=w$ in:
(126) $x\mapsto x-\frac{x.r}{r.r}(1-\lambda)r$
The four mirrors given in the top row generate one of the nine copies of the binary tetrahedral group, acting on the $2$-dimensional subspace perpendicular to $(0,1,-1)$. In particular, this group is a $2$-dimensional complex reflection group. The images of the (anti-Hermitian) Pauli matrices can be written as products of two reflections of order $3$ (with opposite determinants). This representation is equivalent to the sum of the trivial representation and the usual representation on a complex $2$-space.
If we expand the complex $1$-spaces to real $2$-spaces, then the six complex scalar multiples of $12$ complex $3$-vectors become two real scalar multiples of $36$ real $6$-vectors, and the corresponding $36$ real reflections generate the Weyl group of $E_{6}$. In other words, the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ is a subgroup of the Weyl group of $E_{6}$, and is in fact a maximal subgroup of the rotation part of the Weyl group. It is possible that this fact may have led to the appearance of $E_{6}$ symmetry in certain Grand Unified Theories (GUTs) such as [17, 18]. Unfortunately, however, the loss of the unitary structure of the $3$-dimensional representation appears to result in a loss of contact to real world physics. For this reason we are not considering models of $E_{6}$ type in this paper, but sticking rigorously to models that can be justified by experiment.
3. Representations and characters
3.1. Continuous gauge groups
So far we have only looked at the combinatorial structure of the group, which gives a qualitative but not quantitative picture of the underlying physics. In order to introduce measurable physical quantities such as mass, momentum and energy, and allocate them to elementary (or composite) particles we need to introduce real (or complex or quaternionic) representations, and use real (or complex or quaternionic) gauge groups to choose coordinates for the representations.
There are three fundamental representations, described above, given by the embeddings of finite groups in Lie groups as follows:
(127) $Z_{3}$ $\subset U(1)$
(128) $Q_{8}\rtimes Z_{3}$ $\subset Sp(1)\cong SU(2)$
(129) $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ $\subset U(3)$
These embeddings show how the gauge groups $U(1)$, $SU(2)$ and $SU(3)$ of the Standard Model arise from the finite group. In each case, the real scalar factor has been taken out of the gauge group, since it only defines a unit of measurement, rather than a genuine physical property.
In the case of $U(1)$ acting on a complex $1$-dimensional representation, we are left therefore with one physical parameter to measure, which can be represented as an angle, or phase, in the complex plane. In the case of $SU(2)$ acting on a $1$-dimensional quaternionic representation, there are three such angles. Therefore by studying the representations of the binary tetrahedral group $Q_{8}\rtimes Z_{3}$, we might expect to find four of the fundamental mixing angles of the Standard Model, including the electro-weak mixing angle (Weinberg angle). A tentative identification of these four angles is obtained in [9].
In the case of $U(3)$, there are five independent angles in the $3$-dimensional complex representation, bringing the total number of mixing angles up to $9$, enough for the Weinberg angle plus four each in the CKM (Cabibbo–Kobayashi–Maskawa [19, 20]) and PMNS (Pontecorvo–Maki–Nakagawa–Sakata [21, 22]) matrices.
3.2. Clifford theory
The standard method of constructing representations and characters of semi-direct products is Clifford theory [23], which is also used more generally whenever a group has a normal subgroup. Since the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ has $6$ normal subgroups, forming a single chain, we can work our way down the chain, from the largest normal subgroup to the smallest, adding new representations at each stage. The method is standard, and the calculations straightforward, so I do not give many details.
At the first stage, we have only the trivial quotient group, and the trivial representation. At the second stage, we have a quotient $Z_{3}$, which adds a $2$-dimensional real representation, or equivalently a $1$-dimensional complex representation and its complex conjugate. At the third stage, the quotient group is the alternating group $A_{4}$ on four letters, and the new representation is the monomial representation of the group acting by conjugation on the three Pauli matrices, with signs attached.
At the fourth stage, we have the fundamental pseudoreal representation as quaternions, with $Z_{3}$ as the corresponding ‘inertial quotient’, so that the other representations at this stage are obtained by tensoring with the representations of $Z_{3}$. This gives us a pair of complex conjugate two-dimensional representations, or equivalently a real four-dimensional representation. Putting all these stages together gives us the representations of the binary tetrahedral group, which are very well known and have been widely studied. The character table can be found in [24, p.404] and is reproduced in Table 1 for convenience, in a shorthand form in which conjugacy classes that are scalar multiples of each other are grouped together:
Table 1. Character table of the binary tetrahedral group
±$\begin{array}[]{l|cccc}&\pm 1&K&\pm W&\pm W^{2}\cr\hline\cr\mathbf{1}a&1&1&1&1\cr\mathbf{1}b&1&1&v&w\cr\mathbf{1}c&1&1&w&v\cr\mathbf{3}&3&-1&0&0\cr\mathbf{2}a&\pm 2&0&\mp 1&\mp 1\cr\mathbf{2}b&\pm 2&0&\mp v&\mp w\cr\mathbf{2}c&\pm 2&0&\mp w&\mp v\cr\hline\cr\end{array}$
Of particular significance here is that the three faithful representations are obtained by taking the fundamental representation $\mathbf{2}a$ and tensoring with the representations $\mathbf{1}a$, $\mathbf{1}b$ and $\mathbf{1}c$ of $Z_{3}$. This type of pattern is a general feature of Clifford theory, and is a generalisation of the decomposition of representations of direct products of groups as tensor products of representations of the factors. In particular, we have a discrete version of the decomposition of representations of $U(1)\times SU(2)$, but with the ‘mixing’ of $U(1)$ with $SU(2)$ already incorporated into the mathematical structure. A combinatorial version of the mixing occurs already at the group-theoretical level, but a quantitative estimate of the mixing angle can only be obtained by looking in detail at the coordinates of the representations.
3.3. The Hessian group
Coming now to $G_{27}$, whose character table can be found in [24, p.400] we first take a $1$-dimensional complex representation, with inertial quotient $Z_{3}$, and induce up to the whole group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$, to get an $8$-dimensional real representation, together with its tensor products with the representations of $Z_{3}$. The scalars in $G_{27}$ act trivially on these representations, so they are representations of a group $(Z_{3}\times Z_{3})\rtimes Q_{8}\rtimes Z_{3}$ of order $216$, which has been known as the Hessian group since the 1870s.
The table of characters of the complex irreducible representations of the Hessian group is given in Table 2. The vertical lines delineate normal subgroups, so that the first two columns comprise the normal subgroup $G_{27}/Z_{3}\cong Z_{3}\times Z_{3}$, generated by the modified Gell-Mann matrices modulo scalars, and the next two columns comprise the quaternion group $Q_{8}$ generated by the Pauli matrices. The horizontal lines separate cohorts of characters for the various quotient groups. The top two rows contain the centralizer orders and the sizes of the conjugacy classes, respectively.
Table 2. The character table of the Hessian group
$\begin{array}[]{r|rr|rr|rrrrrr|}&216&27&24&4&18&9&6&18&9&6\cr&1&8&9&54&12&24&36&12&24&36\cr\hline\cr\mathbf{1}a&1&1&1&1&1&1&1&1&1&1\cr\mathbf{1}b&1&1&1&1&v&v&v&w&w&w\cr\mathbf{1}c&1&1&1&1&w&w&w&v&v&v\cr\hline\cr\mathbf{3}a&3&3&3&-1&0&0&0&0&0&0\cr\hline\cr\mathbf{2}a&2&2&-2&0&-1&-1&1&-1&-1&1\cr\mathbf{2}b&2&2&-2&0&-v&-v&v&-w&-w&w\cr\mathbf{2}c&2&2&-2&0&-w&-w&w&-v&-v&v\cr\hline\cr\mathbf{8}a&8&-1&0&0&2&-1&0&2&-1&0\cr\mathbf{8}b&8&-1&0&0&2v&-v&0&2w&-w&0\cr\mathbf{8}c&8&-1&0&0&2w&-w&0&2v&-v&0\cr\hline\cr\end{array}$
For some purposes it is useful to replace a pair of complex conjugate characters by their real sum, and to fuse a conjugacy class with its inverse class. This give us a simplified table (see Table 3).
Table 3. Real characters of the Hessian group
$\begin{array}[]{c|rr|rr|rrr|}&1&1&1&1&1&1&1\cr U(1)=SO(2)&2&2&2&2&-1&-1&-1\cr\hline\cr SO(3)&3&3&3&-1&0&0&0\cr\hline\cr SU(2)&2&2&-2&0&-1&-1&1\cr U(2)&4&4&-4&0&1&1&-1\cr\hline\cr SU(3)&8&-1&0&0&2&-1&0\cr U(3)&16&-2&0&0&-2&1&0\end{array}$
The first column gives the name of the compact Lie group generated by the finite group, from which we can see a close relationship to the gauge group of the Standard Model. The even-numbered rows contain the basic building blocks $U(1)$, $SU(2)$ and $SU(3)$ respectively, and the other rows contribute to the ‘mixing’ between these three groups.
From this character table one can easily calculate the anti-symmetric square of $\mathbf{8}a$, and we find
(130) $\Lambda^{2}(\mathbf{8}a)=\mathbf{2}b+\mathbf{2}c+\mathbf{8}a+\mathbf{8}b+\mathbf{8}c.$
In particular, there is a unique anti-symmetric (anti-commutative) algebra structure on $\mathbf{8}a$, namely the Lie algebra. The symmetric square is similarly
(131) $S^{2}(\mathbf{8}a)=\mathbf{1}a+\mathbf{3}a+\mathbf{8}a+\mathbf{8}a+\mathbf{8}b+\mathbf{8}c.$
Hence there are two independent symmetric (commutative) products on $\mathbf{8}a$ as well. One can therefore construct a wide variety of different algebra structures on $\mathbf{8}a$, some of which have been studied in the physical literature [25, 26].
Indeed, the corresponding representations of the compact Lie group $SU(3)$ are as follows:
(132) $\Lambda^{2}(\mathbf{8})$ $=\mathbf{8}+\mathbf{10}+\overline{\mathbf{10}}$
(133) $S^{2}(\mathbf{8})$ $=\mathbf{1}+\mathbf{8}+\mathbf{27}$
so that there is a (unique up to scalars) commutative algebra invariant under $SU(3)$, plus an infinite family of algebras that are neither commutative nor anti-commutative.
Table 4. Abbreviated table of characters for quarks
$\begin{array}[]{r|rr|rr|rrrrrr|}\mathbf{3}b&3&0&-1&-1&t&0&t&-t&0&-t\cr\mathbf{3}c&3&0&-1&-1&vt&0&vt&-wt&0&-wt\cr\mathbf{3}d&3&0&-1&-1&wt&0&wt&-vt&0&-vt\cr\hline\cr\mathbf{9}&9&0&-3&1&0&0&0&0&0&0\cr\hline\cr\mathbf{6}a&6&0&2&0&-t&0&t&t&0&-t\cr\mathbf{6}b&6&0&2&0&-vt&0&vt&wt&0&-wt\cr\mathbf{6}c&6&0&2&0&-wt&0&wt&vt&0&-vt\end{array}$
3.4. The faithful characters
Finally we take a $3$-dimensional complex faithful representation of $G_{27}$, with inertial quotient $Q_{8}\rtimes Z_{3}$, so that Clifford theory implies that we must now tensor with all the representations of the latter group. Let us arrange the conjugacy classes grouped together into scalar multiples. Then we only need to specify the character value on one of these classes, and multiply by the appropriate scalars to get the others. In three cases an element is conjugate to its scalar multiples, so that the character values are all 00, while in the other seven they are not. The $14$ faithful complex characters are those given in Table 4 together with their complex conjugates.
It is not really necessary to write out this table in full, since, as noted above, the seven rows are obtained by multiplying the first row with the seven irreducible characters of the binary tetrahedral group, that is the first seven characters of Table 2. If we take the latter to describe properties of leptons, then Table 4 describes the corresponding properties of quarks.
Indeed, if we have the representations $\mathbf{1}b$, $\mathbf{2}a$ and $\mathbf{3}b$ then we can calculate all the other representations easily as tensor products, subject only to the relations
(134) $\mathbf{2}a\otimes\mathbf{2}a$ $=\mathbf{1}a+\mathbf{3}a$
(135) $\mathbf{3}b\otimes\overline{\mathbf{3}b}$ $=\mathbf{1}a+\mathbf{8}a.$
Hence it is only necessary to gauge the representations $\mathbf{1}b$, $\mathbf{2}a$ and $\mathbf{3}b$, using the gauge groups $U(1)$, $SU(2)$ and $SU(3)$ respectively.
4. Computing gauges
4.1. The electroweak gauge
The $U(1)$ gauge group contains only one parameter, or ‘mixing angle’, that identifies the directions of the real and imaginary parts of the representation. This angle has been computed in [9] as $33.024^{\circ}$. The method of calculation was to superimpose the mass axis on an equilateral triangle marked with the three generations of electrons on the vertices. Hence it is expected that this mixing angle is one of the lepton mixing angles in the Standard Model. However, it should be noted that this is an angle between the mass axis and one of the edges of the equilateral triangle, and it is not clear a priori which edge of the triangle to use, or in which direction. Of the six possible angles, only this one is consistent with the experimental values of the mixing angles. Hence we are not making a prediction here, but fitting the model to the experimental data, using a discrete fitting parameter of $60^{\circ}$.
The $SU(2)$ gauge group similarly contains exactly three independent parameters. The angles calculated in [9] were $28.165516^{\circ}$, $66.7276^{\circ}$ and $2.337325^{\circ}$. The method here was to identify discrete quaternions appropriate for four particles (the three generations of electron, plus a proton) that are ‘fundamental’ as far as electroweak interactions are concerned, and again superimpose a mass axis. There is again a discrete set of possible angles to measure, and again we are fitting the model to the experimental data rather than making a specific prediction.
We should certainly expect one of these angles to be the electroweak mixing angle, but it is not clear a priori whether the other two are lepton-mixing (PMNS) or quark-mixing (CKM) angles. However, if we take into account that the embedding of $Q_{8}\rtimes Z_{3}$ into $SU(3)$ is described by the representation $\mathbf{1}b+\mathbf{2}c$, we should expect one ‘scalar’ (i.e. CP-violating phase) and one generation-mixing angle.
It is therefore proposed to identify these angles, calculated from the directions of the mass axis in the two cases, with four of the Standard Model mixing angles, as follows:
(136) $\theta_{12}^{PMNS}$ $={33.41^{\circ}}^{+.75^{\circ}}_{-.72^{\circ}}$
(137) $\theta_{W}$ $=28.172^{\circ}\pm.021^{\circ}$
(138) $\delta_{CP}^{CKM}$ $=68.8^{\circ}\pm 4.5^{\circ}$
(139) $\theta_{23}^{CKM}$ $=2.38^{\circ}\pm.06^{\circ}.$
4.2. The strong gauge
The $SU(3)$ gauge group acts on a unitary $3$-space, which in principle requires $8$ ‘mixing angles’ to define a unitary basis. But we have already identified $3$ of them from an embedding of $SU(2)$ in $SU(3)$ (the fourth mixing angle above requires an extension to $U(3)$). Hence there are just five more mixing angles to identify.
In the Standard Model these are
(140) $\theta_{12}^{CKM}$ $=13.04^{\circ}\pm.05^{\circ},$
(141) $\theta_{13}^{CKM}$ =.$201^{\circ}\pm.011^{\circ},= .$
(142) $\theta_{13}^{PMNS}$ $={8.54^{\circ}}^{+.11^{\circ}}_{-.12^{\circ}},$
(143) $\theta_{23}^{PMNS}$ $={49.1^{\circ}}^{+1.0^{\circ}}_{-1.3^{\circ}},$
(144) $\delta_{CP}^{PMNS}$ $={197^{\circ}}^{+42^{\circ}}_{-25^{\circ}}.$
In principle it should be possible to calculate these from an appropriate set of particle masses, but it is far from obvious how many masses we need, or what the formulae are. Certainly we need at least six masses, which must presumably include all the quark masses, but it seems more likely that we need to include the three electron masses and the proton mass as well, making ten in all.
4.3. Identifying particles
However it is done, we must start by identifying the particles we use with certain vectors or subspaces in $\mathbb{C}^{3}$. Therefore we must begin by examining the geometry of this $3$-dimensional space, using the complex reflection group as a guide. The most important vectors are those that lie in the $12$ ‘mirrors’. The particular scalar multiple chosen is irrelevant for defining the reflection, so to begin with at least we can ignore the scalar multiplication. The Hessian group then acts on this set of $12$ one-spaces in four triplets, as shown in (125).
The reflections in any mirror fix the mirrors in its column, and cycle the other three columns. The normal subgroup $G_{27}/Z_{3}$ of order $9$ contains four subgroups of order $3$, each of which fixes the mirrors in one column, and cycles the mirrors in the other three columns (not always in the same direction!). It is therefore tempting to identify these $12$ mirrors with the $12$ types of elementary fermions, and use the scalars for unobservable properties such as colours of quarks and spin direction of electrons. However, the details of such an identification are tricky.
An alternative approach might be to fix a particular mirror, say $(t,0,0)$, so that the reflections rotate the mirror itself, to $(vt,0,0)$ and $(wt,0,0)$, fix the two perpendicular mirrors $(0,t,0)$ and $(0,0,t)$, and move the other $9$ mirrors. We can then perhaps identify the last $9$ mirrors as three generations of electrons and up and down quarks in some way, in order to get masses into the picture. Then we might use the mirror itself for a tenth mass such as a proton, or alternatively treat one or all of the three mirrors in the first column as neutrinos. The latter suggestion appears to mix the electron generation with the spin direction for a neutrino, which is exactly the kind of mathematical property that is needed both for the chirality of the weak interaction detected by the Wu experiment [27] and for explaining neutrino oscillations [28].
There is another set of $9$ one-spaces defined as the fixed spaces of the $9$ copies of $Q_{8}\rtimes Z_{3}$. These are the images of $(0,1,-1)$ under coordinate permutations and scalar multiplications on the coordinates. If we fix a particular copy of the group, then the $9$ one-spaces split as $1+8$, which suggests a splitting of the massless bosons in the Standard Model into one photon plus eight gluons. If we have identified the individual coordinates with neutrinos as suggested above, then there is a possible interpretation of gluons (and the photon) as bound states of a neutrino and an anti-neutrino. Indeed, if we take the neutrinos as a basis for $\mathbf{3}b$ and the anti-neutrinos as a basis for $\overline{\mathbf{3}b}$, as above, then the bound states should lie in $\mathbf{3}b\otimes\overline{\mathbf{3}b}=\mathbf{1}a+\mathbf{8}a$.
5. Geometry
5.1. Finite fields
The mathematical structure is best elucidated by regarding $1,v,w$ not as complex numbers, but as abstract elements of the field of order $4$. This is a field of characteristc $2$, so that $2=0$, $-1=1$ and $3=1$. The equation $01+v+w=$ 0 still holds, as in the complex numbers, so that $1+v=w$ and $1+w=v$, as well as $v+w=1$. The multiplicative rules $v^{2}=w$, $w^{2}=v$ and $vw=1$ also hold just as before, and the involution that swaps $v$ with $w$ is the appropriate analogue of complex conjugation. Denote this field by $\mathbb{F}_{4}$, and denote the $n$-dimensional unitary group by $U(n,\mathbb{F}_{4})$.
Then it is easy to see that $U(1,\mathbb{F}_{4})$ is the cyclic group of order $3$, consisting of all non-zero $1\times 1$ matrices. The whole group of order $648$ constructed above is then $U(3,\mathbb{F}_{4})$. But the quaternion group and the binary tetrahedral group do not have obvious names in terms of $\mathbb{F}_{4}$.
The reason for this is that the Pauli matrices are more naturally written over the field $\mathbb{F}_{9}$ of order $9$, whose non-zero elements are ±$\pm 1$, ±$\pm i$ and ±$\pm 1\pm i$. We can then see that the binary tetrahedral group is isomorphic to $SU(2,\mathbb{F}_{9})$. From this point of view, therefore, the reason for the complicated mixing between $SU(2)$ and $SU(3)$ in the Standard Model is ultimately a complicated relationship between the finite fields of orders $4$ and $9$. The particular feature that makes the finite groups behave completely differently from Lie groups is the chain of quotient maps
(145) $U(3,\mathbb{F}_{4})\rightarrow SU(2,\mathbb{F}_{9})\rightarrow U(1,\mathbb{F}_{4})$
This provides a unification of $U(3)$ and $SU(2)$ that is specific to these finite fields, and occurs nowhere else in mathematics. In particular, such unification cannot be achieved with Lie groups over the complex numbers.
The $3$-dimensional space over $\mathbb{F}_{4}$ on which $U(3,\mathbb{F}_{4})$ acts contains $634^{3}-1=$ 63 vectors, each in a set of $3$ scalar multiples, and therefore forming $63/3=21$ one-dimensional subspaces. These $21$ ‘points’ of the geometry divide into $12$ ‘odd’ (non-singular) and $9$ ‘even’ (singular) points. Each point is perpendicular to a ‘line’. The $12$ non-singular lines each contain $2$ non-singular and $3$ singular points, while the $9$ singular lines each contain $1$ singular point and $4$ non-singular points.
The points can be arranged as
(152) $\begin{array}[]{|cccc|}100&111&v11&w11\cr 010&1vw&11v&1w1\cr 001&1wv&1v1&11w\end{array}\qquad\begin{array}[]{|ccc|}011&101&110\cr 0vw&w0v&vw0\cr 0wv&v0w&wv0\end{array}$
to exhibit the symmetry. These are just the reductions to $\mathbb{F}_{4}$ of the $12$ ‘mirrors’ (corresponding to $Z_{3}$ subgroups) and $9$ copies of the Pauli matrices (corresponding to $Q_{8}$ subgroups), so that all of the above discussion of particles applies equally to the finite geometry.
5.2. Interpretations and speculations
In particular, we can interpret the $12$ non-singular points as the fundamental fermions, arranged into four columns of three generations (flavours) each, and it would seem to make sense to interpret the first two columns as leptons, and the last two as quarks, although the mathematics allows us an arbitrary choice here. It is noticeable, however, that in this interpretation there is no single ‘generation symmetry’ that acts on all four columns simultaneously. Instead, there is a group of order $9$ containing elements of order $3$ that act on any three of the four columns.
This is physically reasonable, since it is known that quark generations behave differently from lepton generations in the weak interaction, so that at least two different generation symmetries are required in order to describe what is experimentally observed. We can choose the two generation symmetries in a variety of ways, for example an electron generation symmetry that acts on electrons but not neutrinos, and a neutrino generation symmetry that acts on neutrinos but not electrons. Such a choice shows explicitly that the neutrinos can change flavour independently of the electrons, and therefore models the observed behaviour of neutrino oscillations between the three flavours.
Thus the proposed splitting of the ‘generation’ symmetry into two separate triplet symmetries is supported by experiment, at least to a certain extent. But there is a crucial difference between the electron generation symmetry, in which the mass changes very obviously by several orders of magnitude between generations, and the neutrino generation symmetry, in which no mass can be detected directly at all, but is instead inferred from the details of the theoretical model. It is reasonable to suppose, therefore, that the electron generation symmetry acts to change the mass of the quarks, while the neutrino generation symmetry has no detectable effect on quark mass.
In other words, the neutrino generation symmetry effectively acts on the colours of the quarks rather than their generation. But this implies that the neutrino generation should rather be interpreted as a colour, and that neutrino oscillation should be interpreted as a change in colour, and must therefore be mediated by gluons. This radical proposal has a number of interesting consequences. First, it allows neutrinos to change generation/colour by an interaction (with a massless gluon), rather than spontaneously, and therefore it allows this to happen even if the neutrinos have exactly zero mass. It therefore does not require a hypothesis of non-zero neutrino mass. Second, it allows colours to escape from hadrons, attached to neutrinos, and therefore removes the need to explain ‘colour confinement’. Third, it allows free massless gluons to exist, which are then available as mediators for a quantum theory of gravity. Indeed, it is possible that it even allows bound states of two gluons (with different colour charges) to form spin $2$ gravitons.
6. Spinors
6.1. The Dirac algebra
In the first part of this paper, Sections 1 to 5, we have been concerned only with the gauge group of the Standard Model, and deriving its properties from the properties of a particular complex reflection group of order $648$. We have made no mention of spin or the Dirac matrices [29], which are usually treated more-or-less independently of the gauge groups. However, the formalism of electro-weak mixing in the Standard Model does involve a projection with $1-\gamma^{5}$, which effectively ‘mixes’ part of the gauge group with the Dirac algebra describing spin. That is, although $\gamma^{5}$ commutes with the relativistic spin group generated by even products of $\gamma^{0},\gamma^{1},\gamma^{2}$ and $\gamma^{3}$, as required by the Coleman–Mandula Theorem [30], it does not commute with the odd products.
If we are to integrate the Dirac algebra into the unification scheme proposed above, then we require a non-trivial action of the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$, either on the Dirac algebra itself, or on a finite analogue of the Dirac algebra. The obvious choice is to use the complex representation $\mathbf{1}b+\mathbf{3}b$ for the Dirac spinors, and tensor it with itself or its complex conjugate to get the representation on the Dirac algebra.
However, this does not actually work, for the simple reason that it does not contain a copy of $\mathbf{2}a$, which is required for the weak force to act on leptons. It leads to a model of Pati–Salam [31] type, based on $SU(4)$, but does not unify $SU(4)$ with $SU(2)_{L}$. In fact we need a $4$-dimensional quaternionic representation, based on a complex representation something like $\mathbf{2}a+\mathbf{3}b+\overline{\mathbf{3}b}$. It turns out to be easier to build the algebra first, and then get the group to act on it. However, there is no action of the modified Gell-Mann matrices on the finite group generated by the Dirac matrices as they stand. Therefore it is necessary to extend the Dirac matrices by adding at least one more generator. But it is indeed possible to start with a finite group generated by six (rather than the standard five) Dirac matrices, and get the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ to act on this group.
In fact there is a unique possible action of the proposed finite analogue of the gauge group of the Standard Model on an appropriate finite analogue of the Dirac matrices. This requires extending the Dirac matrices from complex to quaternionic, in order to incorporate three generations of fermions, and extends the usual group of order $64$ to order $128$. The resulting ‘unification’ of Dirac, Gell-Mann and Pauli matrices into a single structure forms a group of order $82944$, consisting of $4\times 4$ quaternion matrices, that is a subgroup of a quaternionic reflection group [32, 33].
If this group is useful for physics, then it certainly goes beyond the Standard Model of Particle Physics, because it allows the Gell-Mann matrices and the Pauli matrices to act on the Dirac matrices, instead of commuting with them. In other words, it allows particle physics to determine the shape of spacetime, and possibly thereby to include a quantum theory of gravity.
6.2. A $4$-dimensional group
The Dirac matrices are $4\times 4$ complex matrices that have a similar structure to that of the Pauli matrices and Gell-Mann matrices. But in this case both Hermitian and anti-Hermitian matrices are required, and the group that is normally used is $SL(2,\mathbb{C})$, rather than $SU(4)$ or $SL(4,\mathbb{R})$. The Dirac matrices are all unitary, however, and they generate a group of order $642^{6}=$ 64. A more natural analogue of $Q_{8}$ and $G_{27}$, however, is a so-called extraspecial group $E_{128}$ of order $128$ obtained by adjoining a complex conjugation operator.
In principle, there is a choice as to whether the complex conjugation operator squares to $+1$ or $-1$, but it turns out that in order to fit the Dirac matrices together with the Gell-Mann matrices it is necessary to take the $-1$ case. Then the representation is quaternionic, which means it can also be written as a group of $4\times 4$ quaternion matrices. The group $E_{128}$ requires six generators, such as
(165) $\begin{pmatrix}i&0&0&0\cr 0&i&0&0\cr 0&0&i&0\cr 0&0&0&i\end{pmatrix},\quad\gamma^{3}=\begin{pmatrix}0&0&1&0\cr 0&0&0&-1\cr-1&0&0&0\cr 0&1&0&0\end{pmatrix},\quad i\gamma^{1}\gamma^{2}\gamma^{3}=\begin{pmatrix}0&0&-1&0\cr 0&0&0&-1\cr 1&0&0&0\cr 0&1&0&0\end{pmatrix},$
(178) $\begin{pmatrix}j&0&0&0\cr 0&j&0&0\cr 0&0&j&0\cr 0&0&0&j\end{pmatrix},\quad\gamma^{1}=\begin{pmatrix}0&0&0&1\cr 0&0&1&0\cr 0&-1&0&0\cr-1&0&0&0\end{pmatrix},\quad i\gamma^{0}\gamma^{2}=\begin{pmatrix}0&0&0&1\cr 0&0&-1&0\cr 0&1&0&0\cr-1&0&0&0\end{pmatrix}.$
These generators are arranged in such a way that the three columns generate three mutually commuting copies of the quaternion group $Q_{8}$.
The names above are given according to the standard Bjorken-Drell convention [34], but many other arrangements are possible. For example, it is possible to choose $\gamma^{1}\gamma^{2}$, $\gamma^{2}\gamma^{3}$ and $\gamma^{1}\gamma^{3}$ as elements of one of the $Q_{8}$ factors, and $i\gamma_{0}$, $i\gamma_{5}$ and $\gamma_{0}\gamma_{5}$ as elements of the other $Q_{8}$ factor. This choice gives a useful mathematical separation of the properties that depend on a direction in space (such as spin and momentum) from those that do not (such as energy, mass and weak isospin).
It is well-known that the outer automorphism group of this group of order $2^{7}$ is isomorphic to the Weyl group of type $E_{6}$, so that it is straightforward to calculate any necessary properties. In fact there are $120$ subgroups isomorphic to $Q_{8}$, and they are all equivalent under the automorphism group, so we can take any one of them to represent the non-relativistic spin group as above. We then need to choose $\gamma^{0}$ and $\gamma^{5}$ from the nine elements of order $2$ (up to sign) that commute with this copy of $Q_{8}$. The automorphism group acts transitively on these $9$ elements, and the stabiliser of one of them acts transitively on the four that anti-commute with it. Hence all choices of Dirac matrices are mathematically equivalent. In particular, there is no reason to choose the Bjorken–Drell convention, or the ‘chiral’ convention (in which $\gamma^{0}$ and $\gamma^{5}$ are swapped, relative to the Bjorken–Drell convention), if other conventions provide simpler notation for physically important concepts. Our conventions, and the reasons behind them, will be explained in Section 6.4.
6.3. A quaternionic reflection group
The action of $G_{27}$ on $E_{128}$ is more difficult to calculate, but much of the work has been done for us in [32, 33], though with some errors in [33] we need to watch out for. It is probably easiest to start from the quaternionic reflection group called $S_{3}$ in [32], which happens to be an extension of $E_{128}$ by the rotation subgroup of the Weyl group of type $E_{6}$. Quaternionic reflections are given by the same formula (126) as for complex reflections, with appropriate choices for the quaternionic scalar $\lambda$. In order to ensure that reflections are linear maps we should check the order of multiplication, since the quaternions are non-commutative. (Note that the conventions in [32, 33] are opposite to each other, and we follow the latter.) The particular order given in (126) ensures that replacing $x$ by a scalar multiple $\mu x$ gives the same scalar multiple on the right hand side. Moreover, replacing $r$ by a scalar multiple $\mu r$ replaces $x.r$ by $(x.r)\bar{\mu}$, so changes the scalar $1-\lambda$ to $1-\mu^{-1}\lambda\mu$. Hence reflections of order $2$, with $\lambda=-1$, are invariant under this operation, but other reflections are not.
The group $G_{27}$ can be generated by the products of pairs of reflections of order $2$ in the mirrors defined by
(179) $(2,0,0,0),(\pm 1,1,1,1),(\pm 1,i,j,k),(\pm 1,j,k,i),(\pm 1,k,i,j)$
Indeed, we can take the products of the first reflection with the other $8$ to represent the $8$ Gell-Mann matrices. Generators for the full group of order $648$ can then be taken as
(192) $\frac{1}{2}\begin{pmatrix}-1&1&1&1\cr-1&1&-1&-1\cr-1&-1&1&-1\cr-1&-1&-1&1\end{pmatrix},\quad\begin{pmatrix}1&0&0&0\cr 0&i&0&0\cr 0&0&j&0\cr 0&0&0&k\end{pmatrix},\quad\begin{pmatrix}1&0&0&0\cr 0&0&1&0\cr 0&0&0&1\cr 0&1&0&0\end{pmatrix}$
The first matrix above represents one of the Gell-Mann matrices in $G_{27}$, the second one of the Pauli matrices in $Q_{8}$, and the last is a generator for $Z_{3}$.
6.4. Dirac matrices
The Dirac matrices for three generations generate the extraspecial group $E_{128}$, which is a commuting product of three copies of $Q_{8}$. There are $40$ different ways in which the group can be written as such a product, but all are equivalent under the action of the Weyl group of $E_{6}$ as the outer automorphism group. Hence we may take the factorisation as the columns in (165) as the ‘standard’ one. However, the appropriate interpretations of the three factors are far from clear.
For example, as a first tentative suggestion, we can separate the quaternions $i,j,k$ to generate one copy of $Q_{8}$ to describe the three generations, and use the real matrices for two other copies of $Q_{8}$ to describe spin and weak isospin (or something related to weak isospin). It therefore makes sense to separate the spin copy of $Q_{8}$ from the ‘weak isospin’ copy, as follows:
(201) $\gamma^{1}\gamma^{2}=\begin{pmatrix}0&0&1&0\cr 0&0&0&-1\cr-1&0&0&0\cr 0&1&0&0\end{pmatrix},\quad\gamma^{0}\gamma^{5}=\begin{pmatrix}0&0&-1&0\cr 0&0&0&-1\cr 1&0&0&0\cr 0&1&0&0\end{pmatrix},$
(210) $\gamma^{2}\gamma^{3}=\begin{pmatrix}0&0&0&1\cr 0&0&1&0\cr 0&-1&0&0\cr-1&0&0&0\end{pmatrix},\quad i\gamma^{0}=\begin{pmatrix}0&0&0&1\cr 0&0&-1&0\cr 0&1&0&0\cr-1&0&0&0\end{pmatrix}.$
This allocation shows that we can only get an energy term in $\gamma^{0}$ if we attach a generation label $i$, $j$ or $k$. Indeed, the same is true for the momentum terms in $\gamma^{1}$, $\gamma^{2}$ and $\gamma^{3}$. To see this, we can take generators for the real Clifford algebra $Cl(3,1)$ as follows:
(219) $i\gamma^{3}=\begin{pmatrix}1&0&0&0\cr 0&-1&0&0\cr 0&0&1&0\cr 0&0&0&-1\end{pmatrix},\quad i\gamma^{2}=\begin{pmatrix}0&0&0&1\cr 0&0&-1&0\cr 0&-1&0&0\cr 1&0&0&0\end{pmatrix},$
(228) $i\gamma^{1}=\begin{pmatrix}0&1&0&0\cr 1&0&0&0\cr 0&0&0&1\cr 0&0&1&0\end{pmatrix},\quad i\gamma^{0}=\begin{pmatrix}0&0&0&1\cr 0&0&-1&0\cr 0&1&0&0\cr-1&0&0&0\end{pmatrix}.$
and extend to the complex Clifford algebra, or $Cl(4,1)$, by adjoining
(233) $\gamma^{5}=\begin{pmatrix}0&i&0&0\cr-i&0&0&0\cr 0&0&0&-i\cr 0&0&i&0\end{pmatrix}$
With our conventions, the matrices $\gamma^{0}$, $\gamma^{1}$, $\gamma^{2}$, $\gamma^{3}$ and $\gamma^{5}$ are all real matrices times $i$. This enables us to reduce from the proposed three-generation model of Dirac matrices to the standard one-generation model by choosing an appropriate imaginary ‘scalar’ $i$. But of course, the symmetry allows us to choose $j$ or $k$ for the other two generations, if this our choice of interpretation. In any case, the choice of real matrices for a model without a generation label (or perhaps better, a model without mass) requires the reversed signature, but only requires $Cl(1,3)$ rather than an extension to a $5$-dimensional ‘spacetime’.
The full $4\times 4$ quaternion matrix algebra can also be interpreted as a Clifford algebra for a $6$-dimensional real space with signature $(6,0)$, $(5,1)$, $(2,4)$ or $(1,5)$. Many models of this type have been proposed [31, 35, 36, 37], including the Pati–Salam model based on $SU(4)=Spin(6)$, and the twistor theory of Penrose, based on the Clifford algebra $Cl(2,4)$ and the corresponding spin group isomorphic to $SU(2,2)$. Twistors can equally well be given the opposite signature, so that the matrix algebra underlying $Cl(4,2)$ is the algebra of $8\times 8$ real matrices. Models of this type have also been proposed [38], but it would appear from this discussion that the signature $Cl(5,1)$ is most likely to provide the closest fit to the Standard Model, and to observed physics, depending of course on the details of the physical interpretation of the mathematical structure.
The Dirac matrices above now act on rows (mathematicians’ convention) or columns (physicists’ convention) of four quaternions, as a generalisation to three generations of the standard Dirac spinors, that are columns of four complex numbers. The projections with $1\pm\gamma^{5}$ onto left-handed and right-handed (Weyl) spinors are a little more complicated than in the standard conventions, and also depend on the generation. But this is a small price to pay for incorporating the generation structure of the elementary fermions directly into the Dirac spinors. At the same time, we see that three generations only require twice as many spinors as one generation, as has also been pointed out in [16, 39]. It is usually assumed that three generations require three times as many spinors as one [40].
6.5. Twistors
The Dirac generators $\gamma^{0}$, $\gamma^{1}$, $\gamma^{2}$, $\gamma^{3}$ and $\gamma^{5}$ generate the Clifford algebra $Cl(2,3)$. Hence in order to obtain $Cl(2,4)$ it is necessary and sufficient to add an extra generator $\gamma^{6}$ that anti-commutes with all of the five generators, and squares to $-1$. In both the Bjorken–Drell convention and the chiral basis the required copy of $\gamma^{6}$ is
(238) $\gamma_{6}=\begin{pmatrix}0&0&0&j\cr 0&0&-j&0\cr 0&-j&0&0\cr j&0&0&0\end{pmatrix}.$
It is therefore possible to use our matrices to work with twistors, without making any changes to the conventional Dirac matrices.
But if we want to follow Woit’s approach [38], in which a gauge group $SU(2)_{R}$ for quantum gravity appears, then it will be useful to make the factorisation of the group $E_{128}$ into three copies of $Q_{8}$ more prominent. The scalar $i$ appears as the volume element (pseudoscalar) of $Cl(2,3)$, so the simplest approach is to define $\gamma^{6}=j$, which is equivalent to choosing $i\gamma^{0}$, $i\gamma^{1}$, $i\gamma^{2}$ and $i\gamma^{3}$ (and therefore also $i\gamma^{5}$) to be real matrices. A possible choice of such real matrices is given in (219), but there are many others. The choice made here includes (a) a choice of $i\gamma^{3}$ to be diagonal, in order to measure spin in the $z$ direction, as is traditional, and (b) a choice of the coordinate permutation part of $i\gamma^{0}$ that makes $\gamma^{0}\gamma^{2}$ diagonal.
The twistors themselves are much the same as the quaternionic Dirac spinors, and can be represented as quaternionic $4$-vectors, either as rows or as columns according to convention. However, the quaternionic structure is not used in most approaches to twistors, which are considered to lie in spinor representations of $Spin(2,4)\cong SU(2,2)$, and are therefore complex $4$-vectors. Each quaternionic Dirac spinor therefore consists of a chiral pair of Penrose twistors.
7. The structure of the combined group
7.1. Gell-Mann matrices
The full set of $8$ Gell-Mann matrices can be taken as the following four matrices and their conjugate transposes:
(247) $\frac{1}{2}\begin{pmatrix}-1&1&1&1\cr-1&1&-1&-1\cr-1&-1&1&-1\cr-1&-1&-1&1\end{pmatrix},\quad\frac{1}{2}\begin{pmatrix}-1&i&j&k\cr i&1&k&-j\cr j&-k&1&i\cr k&j&-i&1\end{pmatrix},$
(256) $\frac{1}{2}\begin{pmatrix}-1&j&k&i\cr j&1&i&-k\cr k&-i&1&j\cr i&k&-j&1\end{pmatrix},\quad\frac{1}{2}\begin{pmatrix}-1&k&i&j\cr k&1&j&-i\cr i&-j&1&k\cr j&i&-k&1\end{pmatrix}.$
The two real matrices here correspond to the two colourless gluons, and the other six to the coloured gluons. Writing $v=(-1+i+j+k)/2$, the group of order $27$ generated by these matrices contains a centre of order $3$, generated by
(261) $\begin{pmatrix}v&0&0&0\cr 0&0&v&0\cr 0&0&0&v\cr 0&v&0&0\end{pmatrix}.$
This matrix acts on the Dirac matrices to permute the three generations defined by $i$, $j$ and $k$. At the same time, it permutes the three directions of spin, defined by $\gamma^{1}\gamma^{2}$, $\gamma^{2}\gamma^{3}$ and $\gamma^{3}\gamma^{1}$. But it is important to note that replacing $v$ by $w$ in (261) can only be done if the coordinate permutation is also inverted. The chirality of the weak interaction ultimately arises from this important fact, as we shall see in the next section. It is also worth noting that this matrix fixes the quaternionic one-space spanned by $(0,1,v,w)$, on which all the Gell-Mann matrices act trivially.
7.2. Pauli matrices
The (anti-Hermitian) Pauli matrices generating $Q_{8}$ are represented by the diagonal matrices
(274) $\begin{pmatrix}1&0&0&0\cr 0&i&0&0\cr 0&0&j&0\cr 0&0&0&k\end{pmatrix},\quad\begin{pmatrix}1&0&0&0\cr 0&j&0&0\cr 0&0&k&0\cr 0&0&0&i\end{pmatrix},\quad\begin{pmatrix}1&0&0&0\cr 0&k&0&0\cr 0&0&i&0\cr 0&0&0&j\end{pmatrix},$
acting on the spinors. These matrices commute with (261), but are permuted by both the ‘scalar’ $v$ and the coordinate permutation separately. The scalar acts by permuting the three generations, but leaving the spin direction alone, while the permutation acts on the spin direction but not on the generation. The fact that the matrix (261) couples the cyclic ordering of the three generations to the cyclic ordering of the three directions of spin is what is known as the chirality of the weak interaction. Conventionally this chirality is left-handed, so that the copy of $SU(2)$ generated by the Pauli matrices above is denoted $SU(2)_{L}$.
The way chirality is implemented in the Standard Model is by restricting to the third component of weak isospin, represented say by the first matrix in (274), and realising $j$ as complex conjugation, so that the bottom half of the Dirac spinor behaves as a Weyl spinor that is the complex conjugate of the top Weyl spinor. Thus the proposed model is a straightforward generalisation of the Standard Model at this point, and incorporates all three components of weak isospin, not just one.
We can now extend from $Q_{8}$ to the binary tetrahedral group by adjoining any one of the three matrices
(287) $\begin{pmatrix}1&0&0&0\cr 0&0&1&0\cr 0&0&0&1\cr 0&1&0&0\end{pmatrix},\quad\begin{pmatrix}v&0&0&0\cr 0&v&0&0\cr 0&0&v&0\cr 0&0&0&v\end{pmatrix},\quad\begin{pmatrix}v&0&0&0\cr 0&0&0&v\cr 0&v&0&0\cr 0&0&v&0\end{pmatrix}.$
As noted above, the first of these acts on the colours, and on the direction of spin, but not on the generations. The second one acts on the colours and the generations, but not on the direction of spin. The third one acts on all three. We may therefore wish to keep this choice open until such time as it becomes clearer what the implications of this choice might be.
8. Action on spinors and Dirac matrices
8.1. Factorisation of the Dirac matrices
As noted in Section 6.2, the quaternionic Dirac matrices factorise in $40$ different ways as a commuting product of three copies of $Q_{8}$. I have tentatively proposed to identify one copy as a generation or flavour group, and another as a (non-relativistic) spin group. The third therefore has to deal with both relativistic corrections and weak isospin. This interpretation may or may not be suitable: it all depends on how the Gell-Mann and Pauli matrices act on the Dirac matrices in this model.
First note that the Gell-Mann matrices preserve this decomposition, but the Pauli matrices do not. This means that the Pauli matrices (representing the weak interaction) mix the flavour group with the spin group, in particular. This is the characteristic experimental property of beta decay, as measured by the Wu experiment [27] in 1957: the first-generation electron/neutrino flavour is coupled to the direction of spin and the direction of momentum. Hence this mathematical property of the model is fundamentally linked to the ‘chirality’ of the weak interaction, in some way which may repay more detailed investigation.
To illustrate this in more detail, consider the first of the Pauli matrices represented in (274), acting by conjugation on the quaternionic Dirac matrices in (165), but with the labelling of the real part given in (201) and (219). The first row of (165) then maps to
(300) $\begin{pmatrix}i&0&0&0\cr 0&i&0&0\cr 0&0&-i&0\cr 0&0&0&-i\end{pmatrix},\quad\begin{pmatrix}0&0&j&0\cr 0&0&0&j\cr j&0&0&0\cr 0&j&0&0\end{pmatrix},\quad\begin{pmatrix}0&0&-j&0\cr 0&0&0&j\cr-j&0&0&0\cr 0&j&0&0\end{pmatrix}.$
In particular, the ‘flavour’ matrix has been multiplied by the diagonal matrix $\gamma^{0}\gamma^{2}$, which combines elements from both the ‘spin’ and ‘weak isospin’ copies of $Q_{8}$, as we would expect from experiment. Correspondingly, the ‘spin’ matrix has acquired a factor of $j$, corresponding to complex conjugation in the Standard Model, that swaps the left-handed and right-handed Weyl spinors. The ‘weak isospin’ matrix has also acquired a factor of $j$, but how this should be interpreted I am not sure. Indeed, it is probably not appropriate to use the words ‘weak isospin’ for this copy of $Q_{8}$, whose main purpose seems to be to unify kinetic energy in $i\gamma^{0}$ with other types of energy in $i\gamma^{5}$ and $\gamma^{0}\gamma^{5}$.
8.2. Action of Gell-Mann matrices
We now consider the Gell-Mann matrices (247). The first pair, that is the real matrices, representing colourless gluons, commute with the ‘flavour’ $Q_{8}$, consisting of ‘scalar’ matrices. Since they have order $3$, they cannot swap the other two factors, so they fix both the ‘spin’ and ‘weak isospin’ copies of $Q_{8}$. The precise action on spin can be calculated as follows:
(313) $\frac{1}{2}\begin{pmatrix}-1&-1&-1&-1\cr 1&1&-1&-1\cr 1&-1&1&-1\cr 1&-1&-1&1\end{pmatrix}\begin{pmatrix}0&0&1&0\cr 0&0&0&-1\cr-1&0&0&0\cr 0&1&0&0\end{pmatrix}$ $\frac{1}{2}\begin{pmatrix}-1&1&1&1\cr-1&1&-1&-1\cr-1&-1&1&-1\cr-1&-1&-1&1\end{pmatrix}$
(318) $=\begin{pmatrix}0&0&0&1\cr 0&0&1&0\cr 0&-1&0&0\cr-1&0&0&0\end{pmatrix}$
which shows that these gluons have the effect of changing the direction of spin from the $z$ direction to the $x$ and $y$ directions.
The action on the third copy of $Q_{8}$ is similar:
(331) $\frac{1}{2}\begin{pmatrix}-1&-1&-1&-1\cr 1&1&-1&-1\cr 1&-1&1&-1\cr 1&-1&-1&1\end{pmatrix}\begin{pmatrix}0&0&-1&0\cr 0&0&0&-1\cr 1&0&0&0\cr 0&1&0&0\end{pmatrix}$ $\frac{1}{2}\begin{pmatrix}-1&1&1&1\cr-1&1&-1&-1\cr-1&-1&1&-1\cr-1&-1&-1&1\end{pmatrix}$
(336) $=\begin{pmatrix}0&-1&0&0\cr 1&0&0&0\cr 0&0&0&1\cr 0&0&-1&0\end{pmatrix}$
Here we see a conversion from $\gamma^{0}\gamma^{5}$ to $i\gamma^{5}$ (and also to $i\gamma^{0}$), which indicates a mixing of some kind with the weak interaction, dependent on energy.
8.3. Colour
The action of the coloured gluons is a good deal more complicated, since they act by permuting the three copies of $Q_{8}$. Thus they map spin symmetries to flavour symmetries and weak isospin symmetries, which seems to imply that the entire structure of spacetime breaks down once we look at the internal structure of a baryon. In other words, at distances less than about $10^{-15}$m, it no longer makes sense to talk about ‘space’ at all. Dividing by the speed of light to convert distance to time, we obtain a cutoff of around $3\times 10^{-24}$s below which it does not make sense to talk about ‘time’. This is roughly half the lifetime of the most unstable mesons and baryons in the Standard Model.
Note however that the lifetime of the $W$ and $Z$ bosons is an order of magnitude smaller, indicating that the process of beta decay of a neutron takes place on this even smaller scale of both time and space, where spacetime may behave very strangely compared to our expectations based on macroscopic physics.
It is straightforward to compute the action of the coloured gluons on the Dirac matrices, and verify that the action is compatible with (72). Hence the latter notation provides a simpler way to see the action, if we interpret the three complex coordinates as representing the three $Q_{8}$ factors. In this way the diagonal matrices represent the colourless gluons, and the coloured gluons permute the three factors.
8.4. Action on spinors
The above analysis shows that the finite group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ takes the ‘spin’ copy of $Q_{8}$ to a total of $12$ distinct copies within the quaternionic Dirac matrices. Moreover, in each case there is still a choice of $9$ elements to represent the energy in $\gamma^{0}$. In other words, in this model the nuclear forces act to change the shape of spacetime, at least locally. In principle this gives us a mechanism for relating the mass that is created or destroyed by the weak and strong forces (and potentially also the Higgs field) to the shape of spacetime that determines the gravitational field in General Relativity. However, an analysis of the Dirac matrices from this point of view will be enormously complicated, so I make no attempt to do this here.
Indeed, before considering such action on the quaternionic Dirac matrices we should investigate in much more detail the action on the quaternionic Dirac spinors, which is of course quite distinct. Whilst the Standard Model splits the complex Dirac spinor into two pieces, the left-handed and right-handed Weyl spinors, the splitting here is into four pieces, which do not naturally have a complex (Weyl spinor) structure, but rather a quaternionic structure. We can recover the standard description by choosing a complex structure, which breaks the symmetry between the three generations, and at the same time defines a particular copy of the Lorentz group among the many that are available.
The action of the whole group on spinors is easily seen to afford the representation $\mathbf{2}a+\mathbf{3}b+\overline{\mathbf{3}b}$, as was previously suggested. It is easy to calculate that the Gell-Mann matrices act trivially on the quaternionic one-space spanned by $(0,1,v,w)$, and the Pauli matrices act trivially on the quaternionic one-space spanned by $(1,0,0,0)$. The latter therefore replaces the ‘right-handed’ Weyl spinor in the Standard Model, but it is not a Weyl spinor with respect to our choice of Dirac matrices. Possibly this indicates that we need a different choice for the Dirac matrices, changing the labels on the diagonal matrices from $i\gamma^{3}$ and $\gamma^{0}\gamma^{2}$ to something else. However, the model actually contains $9$ copies of the Pauli matrices, which suggests that the real situation is much more complicated than a simple splitting into left-handed and right-handed Weyl spinors.
9. Intimations of quantum gravity
9.1. Choosing a gauge group
In the cases of the weak and strong forces, the gauge group is obtained by exponentiating every non-trivial element of the finite group modulo scalars, that is every non-trivial element of the groups
(337) $Q_{8}/Z_{2}$ $=Z_{2}\times Z_{2},$
(338) $G_{27}/Z_{3}$ $=Z_{3}\times Z_{3}.$
If we adopt the same principle here, we obtain a gauge group of dimension $63$, isomorphic to $SL(4,\mathbb{H})$, from the quotient
(339) $E_{128}/Z_{2}=Z_{2}\times Z_{2}\times Z_{2}\times Z_{2}\times Z_{2}\times Z_{2}.$
In particular, this gauge group is not compact, because the finite group contains elements that square to $+1$ as well as elements that square to $-1$. If we want, we can restrict to the compact part, that is $Sp(4)$, of dimension $36$. However, these gauge groups are not Lorentz-invariant, and indeed $SL(4,\mathbb{H})$ actually contains the Lorentz group.
A better approach is probably to utilise the splitting of $E_{128}$ as a commuting product of three copies of $Q_{8}$, and only exponentiate the elements that lie in one (or perhaps two) of the factors. Indeed, two copies of $Q_{8}$ give a $15$-dimensional gauge group $SL(4,\mathbb{R})$, containing the Lorentz group $SL(2,\mathbb{C})$, and with compact part $SO(4)$. The latter splits into two commuting copies of $SU(2)$, that are sometimes called left-handed and right-handed [38]. It is plausible that these groups can be used to reproduce General Relativity, in which case the third copy of $Q_{8}$ provides a third gauge group $SU(2)$ that is not only Lorentz-invariant, but generally invariant. This is the copy generated by the quaternionic scalars $i$ and $j$.
This copy of $SU(2)$ is not in the Standard Model, and is not in General Relativity. It therefore represents the (only) addition that is required in order to unify particle physics and gravity into a single theory. As a gauge group, it gauges three quantities that link particle physics to gravity. Since the generation symmetry acts on these three quantities, it is reasonable to suppose that they can be taken as the masses of the three generations of electron, relative to a standard unit of mass, say the proton or the neutron. What it must do, therefore, is to reconcile (or calibrate) the inertial masses that are measured in particle physics with the (active) gravitational masses that are measured with Cavendish-type experiments [41, 42, 43, 44].
A similar mathematical idea is proposed in [45], where a group $Spin(3,3)\cong SL(4,\mathbb{R})$ is used to incorporate three generations of fermions by extending the concept of ‘time’ to three dimensions. The difference really is in the interpretation, and in particular the question of which mathematical concept do we label with the physical name of ‘spacetime’. For compatibility with General Relativity, we need to label the $4$-dimensional representation of $SL(4,\mathbb{R})$ as spacetime, and not the $6$-dimensional representation. Of course, this labelling is inconsistent with the conventional identification of $SL(2,\mathbb{C})$ as the Lorentz group $SO(3,1)$, or more precisely the connected component of the identity $SO(3,1)^{o}$, acting on spacetime. But it is required for general covariance of any model of particle physics. Indeed, the real part of the quaternionic $4$-space is the tensor product of spinor representations of two copies of $SU(2)$, so it is most naturally interpreted as a vector, and not a spinor. The subgroups $SL(2,\mathbb{C})$ and $SO(3,1)$ of $SL(4,\mathbb{R})$ are in fact disjoint, and should not be confused with each other.
9.2. Gravi-weak mixing
If we do take the copy of $SU(2)$ generated by the scalar matrices $i$, $j$ and $k$ as the gauge group $SU(2)_{G}$ for quantum gravity, as suggested above, then there is some interesting mixing with the weak copy $SU(2)_{W}$, generated by the diagonal matrices with entries $(1,i,j,k)$, $(1,j,k,i)$ and $(1,k,i,j)$. These two groups do not commute, and the finite group generated by the two copies of $Q_{8}$ in fact has order $2^{8}$. The commutator subgroup consists of all sign changes on the four quaternion coordinates, and modulo signs we have four scalars plus $12$ even permutations of $(1,i,j,k)$.
This group therefore distinguishes four types of ‘spin’, corresponding to the four quaternionic coordinates. In a gravitational context, these four coordinates are labelled by the four coordinates of spacetime, most reasonably in the order $(t,x,y,z)$. Then the $t$-spinor is acted on by gravity but not the weak force, so is a ‘right-handed’ spinor in usual parlance. The $x$-, $y$- and $z$-spinors are then all ‘left-handed’ spinors, acted on by both gravity and the weak force. The allocation to space directions $x$, $y$ and $z$ is in some sense arbitrary, but should preferably be chosen with the ambient gravitational field in mind.
The non-commuting of $SU(2)_{G}$ and $SU(2)_{W}$ implies that the gravitational eigenstates of particles are not compatible with the weak eigenstates. Gravitational eigenstates are usually called ‘mass’ eigenstates, and weak eigenstates are ‘flavour’ eigenstates. In the case of neutrinos, the incompatibility of ‘mass’ and ‘flavour’ eigenstates is well-known, and is the standard explanation for neutrino oscillations. Thus this model includes neutrino oscillations as standard (unlike the Standard Model, which had to add them in later).
In the case of quarks, it is known that there is an incompatibility between the weak and strong eigenstates. It makes sense to regard the strong eigenstates as mass eigenstates, so that we again see a discrepancy between mass eigenstates and flavour eigenstates. In the case of electrons, no such discrepancy exists in the Standard Model, but it does in the proposed model. In this case, we must interpret the weak eigenstates as ‘inertial mass’ eigenstates, since the masses are measured by electromagnetic means, and interpret the gravitational eigenstates as ‘(active) gravitational mass’ eigenstates. Thus the model predicts that inertial mass and active gravitational mass are not the same concept. In other words, it predicts failure of the (strong form of the) Weak Equivalence Principle. Note however that this does not imply the inequivalence of inertial mass and passive gravitational mass. This weaker form of the Weak Equivalence Principle is called ‘universality of freefall’, and is strongly supported by experiment, but does not contradict the proposed model. Unfortunately it is not always easy to tell in the literature which form of the Weak Equivalence Principle is being used in any particular context.
9.3. Speculation
The hope is, therefore, that by not breaking the symmetry of the quaternions, we can also avoid the necessity to choose a definition of inertial frame. Or, to put it another way, the finite group implies that there is a ‘mixing’ between the generation symmetry and the choice of inertial frame. This type of mixing between particle physics and gravity was first considered by Einstein [46] in 1919, but has not found favour in mainstream physics, in which the separation between gravity and particle physics has been rigorously maintained. Nevertheless, the conventional choice of the laboratory frame of reference as ‘inertial’ is fraught with problems, particularly in the case of ‘long-baseline’ experiments spanning two different laboratories, with very different definitions of inertial frame.
The model proposed here implies that such a change in inertial frame should be reflected in a change in ‘generation’ of fermions. Of course, there is no experimental evidence that the generation of an electron is affected in this way, but there is some experimental evidence that the generation (flavour) of a neutrino is [28]. Indeed, it is shown in [9] that by inserting the masses of the three generations of electron into the representations of the $Z_{3}$ quotient of our finite group, the experimental value of the mixing angle between electron neutrino and muon neutrino can be obtained. In other words, the model implies that the electron and neutrino generations are coupled together in a more complicated way than in the Standard Model.
As regards quarks, the evidence is more equivocal, and would require more detailed analysis. But again, the model we are considering here allows us to put the masses of three generations of electron into a quaternion representation of $Q_{8}\rtimes Z_{3}$, together with a proton or neutron, from which the correct experimental value for the mixing angle between second and third generation quarks can be calculated [9]. Again, this suggests a more complicated relationship between lepton and quark generations than is implied by the Standard Model.
9.4. Mass parameters
We have a further $15$ parameters in the coordinates of the $4$-dimensional quaternionic representation on the spinors. If we take at face value the proposed use of this representation for a theory of gravity, then we should expect these parameters to be masses of elementary particles. The Standard Model nowadays has $15$ mass parameters, including $12$ for the elementary fermions, plus the $W$, $Z$ and Higgs bosons.
Finally, note that by restricting the $4$-dimensional representation to subgroups $Z_{3}$, $Q_{8}\rtimes Z_{3}$ and $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ we obtain the first three representations again. It is therefore entirely possible that the $9$ mixing angles of the Standard Model are redundant parameters, and can be derived from the $15$ masses. This is known to be the case for the electro-weak mixing angle, which is derived from the mass ratio of the $W$ and $Z$ bosons. In [9] three more such derivations are proposed.
10. Conclusion
10.1. Summary
The suggestion that physics might be fundamentally discrete, and that the apparently continuous behaviour we often observe is an illusion, goes back a long way. In the context of quantum mechanics, this suggestion was made by Einstein [47] as early as 1935, but no discrete theory was found to rival the continuous models based on the Schrödinger equation and then the Dirac equation. The Standard Model of Particle Physics, largely completed by the mid 1970s, is entirely based on continuous groups, so that the emergence of discrete experimental observations is still not explained in a totally convincing way [48, 49, 50].
Theories beyond the Standard Model are almost without exception continuous theories. A few lone voices, such as ’t Hooft [51], still speak out in favour of Einstein’s dream of a discrete foundation, even if this has to be at the Planck scale. But a convincing discrete model is still lacking, and few people are seriously looking for one. The ones I have put forward in [9, 52, 53, 54, 10, 55] all suffer from drawbacks of various kinds, most notably the lack of any convincing analogue of the colour $SU(3)$, which is an essential and very successful part of the Standard Model. Nevertheless, it is a mathematical fact that finite groups are much more complicated than Lie groups, and quite small finite groups can model properties that cannot be modelled with Lie groups at all.
In this paper, therefore, I have tried to remedy this defect by explicitly proposing a finite analogue or precursor of $SU(3)$. At a combinatorial (qualitative) level it has many desirable properties that mirror known experimental properties of elementary particles, in particular the chirality of the weak interaction, the mixing of electromagnetism and the weak and strong nuclear forces, and neutrino oscillations. Quantitative analysis of masses, mixing angles and probability amplitudes is, however, beyond the scope of this paper, as it requires detailed investigations of coordinates for the representations, that may even (counter-intuitively) depend on the large-scale gravitational environment.
10.2. Further work
The proposed unification of Gell-Mann and Pauli matrices using a 3-dimensional complex reflection group really does little more than provide a discrete foundation for the Standard Model of Particle Physics as it currently stands. Nevertheless, it permits new methods of calculation, using finite geometries, which have been shown to be effective in calculating four of the mixing angles from fundamental mass parameters. A priority for further work is therefore to try to find similar calculations for the remaining five mixing angles.
Further unification with the Dirac matrices cannot be done in the Standard Model, since it is impossible to arrange for the Lorentz group to commute with the whole of the gauge group in this case. Nevertheless, the non-commutation is very subtle, and only really affects properties of neutrinos and gluons, as well as the Higgs field, so that any differences from a commuting theory are hard to detect. What this means is that a unification with Dirac matrices provides a possible mechanism for particle physics to influence the shape of spacetime, and hence couple to gravity. In other words, the proposed unification using a quaternionic reflection group provides a potential foundation for a quantum theory of gravity.
Investigating this possibility further is a priority for future work. The appropriate compact gauge group here is $Sp(4)$, of dimension $36$, but it is also possible to gauge the theory with the noncompact group $SL(4,\mathbb{H})$ of dimension $63$. In this case, there is a subgroup $SL(4,\mathbb{R})$ that might be used for general covariance, in order to reduce any such putative quantum theory of gravity to General Relativity in the appropriate limit. Since $SL(4,\mathbb{R})$ contains no quaternions, it cannot model the three generations of electrons, which suggests that it is the (quantum and/or classical) gravitational properties of electrons [56] that need to be investigated in case of any apparent deviations [57, 58] of gravity from General Relativity.
If a quantum theory of gravity along these lines turns out to be viable, then it might lead to methods for calculating the $15$ mass parameters of the Standard Model from something more fundamental. It is not at all clear what could be ‘more fundamental’, or how it could be used, so that this is at present a highly speculative idea. Nevertheless, it is clear that it will be necessary to examine closely the theoretical and practical definitions of inertial frames of reference in different contexts, and to calculate the effects of transformations between them.
Since $SL(4,\mathbb{H})$ is a real form of $SU(8)$, of Lie type $A_{7}$, and since $A_{7}$ is contained in both $E_{7}$ and $D_{8}$, there are at least two ways to embed this gauge group in (complex) $E_{8}$. It may even be possible to embed it in a real form of $E_{8}$. In the $E_{7}$ case, there are no spinor representations of $SL(4,\mathbb{H})$ in $E_{8}$, which seems to rule this one out [59, 60]. In the $D_{8}$ case, the spinors split into $4+28$ quaternionic dimensions, which may be viable as a splitting into neutrinos versus the charged fundamental fermions. Even so, the model would still be quite unlike the various $E_{8}$ models [16, 45, 61, 62] that have been proposed over a number of years.